### v1

# Verification of Linear Dynamical Systems via O-Minimality of Real Numbers

Abstract.
We study decidability of the following problems for both discrete-time and continuous-time linear dynamical systems. Suppose we are given a matrix $M$, a set $S$ of starting points, and a set $T$ of unsafe points.
  • (a)
    Does there exist $\varepsilon>0$ such that the trajectory of every point in the $\varepsilon$-ball around $S$ avoids $T$?
  • (b)
    Does there exist $\varepsilon>0$ such that every $\varepsilon$-pseudo-orbit of a point $s\in S$ avoids $T$?
These two problems correspond to two different notions of robust safety for linear dynamical systems. Restricting $S$ to be bounded in both questions, and $M$ to be diagonalisable in (b), we prove decidability for discrete-time systems and conditional decidability assuming Schanuel’s conjecture for continuous-time systems. Our main technical tool is the o-minimality of real numbers equipped with arithmetic operations and exponentiation.
1. Introduction
Linear dynamical systems (LDS) are mathematical models widely used in engineering and science to describe systems that evolve over time. A continuous-time LDS in dimension $d$ is given by the differential equation $x^{\prime}(t)=Mx(t)$, where $M\in\mathbb{R}^{d\times d}$ and $x(t)\in\mathbb{R}^{d}$ for $t\geq 0$. The configuration of this system at time $t\geq 0$ satisfies $e^{Mt}\cdot x(0)$. A discrete-time LDS, on the other hand, evolves according to the relation $x(n+1)=Mx(n)$, where $M\in\mathbb{R}^{d\times d}$ and $x(n)\in\mathbb{R}^{d}$ for $n\in\mathbb{N}$. In this model, time is viewed as being discrete, and the trajectory of a point $s\in\mathbb{R}^{d}$ is the sequence $\langle s,Ms,M^{2}s,\ldots\rangle$. Discrete-time LDS arise in program verification in the form of linear loops, which are program fragments of the form
initialise $x\in S$
while $\textrm{while }x\notin T\>\textrm{ do }\>x=Mx$
with $S,T\subseteq\mathbb{R}^{d}$. The Termination Problem for linear loops, in its full generality, is to decide, given $M,S,T$ as above, whether the loop terminates for all $x\in S$ [^1] See (tiwari2004termination, ; braverman2006termination, ; ouaknine2014termination, ) which consider the Termination Problem with the initial sets $\mathbb{R}^{d},\mathbb{Q}^{d}$, and $\mathbb{Z}^{d}$, respectively. In the language of LDS, the Termination Problem asks to decide whether for all $s\in S$, the trajectory of $s$ under $M$ reaches $T$. In this work we consider (variations of) the dual Safety Problem: Given $\langle M,S,T\rangle$ as above, decide whether for all $x\in S$ the trajectory of $s$ under $M$ avoids $T$. This problem arises, for example, when designing systems that are intended to run forever without ever entering an unsafe configuration.
The decidability of the Safety Problem with singleton $S$ has been extensively studied. [^2] The results mentioned in this paragraph hold assuming $M$ has real algebraic entries and $S,T$ are semialgebraic. When discussing decidability, we have to restrict inputs to some class that can be effectively represented and manipulated. For discrete-time LDS and (linear) hyperplane targets, the problem is Turing-equivalent to the famously unsolved Skolem Problem about linear recurrence sequences. Consequently, it is decidable for LDS of dimension at most $4$, and open in dimensions 5 and higher (chonev2016complexity, ). Independently of this Skolem-hardness, it is  known that solving the Safety Problem for singleton initial sets and polytopic (respectively, halfspace) target sets in dimension $4$ (respectively, dimension 5) would lead to major mathematical breakthroughs in Diophantine approximation (chonev2014polyhedron, ; ouaknine2014positivity, ). For continuous-time LDS, the Safety Problem has been studied in (chonev2023zeros, ). They show that for singleton $S$ and hyperplane targets, the Safety Problem is decidable in dimension at most 8 assuming Schanuel’s conjecture, and hard with respect to certain open problems in Diophantine approximation in dimensions 9 and higher. Schanuel’s conjecture (waldschmidt2013diophantine, , Chap. 1.4) is a unifying conjecture in transcendental number theory that yields decidability of the Time-Bounded Safety Problem: Given semialgebraic sets $S,T$, a time bound $N$, and an update matrix $M$, decide whether $e^{Mt}s\notin T$ for all $s\in S$ and $t\in[0,N]$. [^3] Also see Lemma 2 of this paper. The decision procedure of (chonev2023zeros, ) works by (unconditionally) reducing the Safety Problem for singleton $S$ and hyperplane $T$ in dimension at most $8$ to the Time-Bounded Safety Problem.
We will study two modifications of the Safety Problem, motivated by the consideration that systems in practice are often subject to external perturbations. For $S\subseteq\mathbb{R}^{d}$ and $\varepsilon>0$, denote by $B(S,\varepsilon)$ the set $\{x\in\mathbb{R}^{d}\mid\exists s\in S\colon\|x-s\|<\varepsilon\}$. Our first problem is the Robust Safety Problem: Given $M\in\mathbb{R}^{d\times d}$ with real algebraic entries and semialgebraic $S,T\subseteq\mathbb{R}^{d}$, decide whether there exists $\varepsilon>0$ such that the trajectory of every $x\in B(S,\varepsilon)$ in the dynamical system defined by $M$ avoids $T$. Intuitively, in the setting of the Robust Safety Problem we are uncertain about the initial configuration of the system. The class of semialgebraic sets is highly expressive, and includes, among others, Boolean combinations of polytopes and solution sets of polynomial equations. Restricting $S$ to be bounded, we prove conditional decidability for continuous-time systems and full decidability for discrete-time systems. This is in stark contrast to the Safety Problem, which is open and believed to be hard already for singleton $S$ as discussed above.
Theorem 1.
The Robust Safety Problem for continuous-time LDS is decidable for bounded $S$ assuming Schanuel’s conjecture. Moreover, given a positive instance $\langle M,S,T\rangle$ with bounded $S$, we can effectively compute $\tilde{\varepsilon}\in\mathbb{Q}_{>0}$ such that for all $x\in B(S,\tilde{\varepsilon})$ the trajectory of $x$ avoids $T$.
Theorem 2.
The Robust Safety Problem for discrete-time LDS is decidable for bounded $S$. Moreover, given a positive instance $\langle M,S,T\rangle$ with bounded $S$, we can effectively compute $\tilde{\varepsilon}\in\mathbb{Q}_{>0}$ such that for all $x\in B(S,\tilde{\varepsilon})$ the trajectory of $x$ avoids $T$.
To solve the Robust Safety Problem, we introduce the decomposition method. Let us focus on continuous-time systems first. We have to decide whether there exists $\varepsilon>0$ such that the sequence [^4] Here we view the sequence as being indexed over $\mathbb{R}_{>0}$. of sets $(e^{Mt}\cdot B(S,\varepsilon))_{t\geq 0}$ avoids $T$. We first write $e^{Mt}$ as $e^{Ct}e^{Dt}$, where for all $t\geq 0$, $e^{Ct}$ is a scaling matrix, meaning that its eigenvalues are all non-negative reals, and $e^{Dt}$ is a rotation matrix, meaning that it is diagonalisable and its eigenvalues all lie on the unit circle in $\mathbb{C}$. The decision problem can then be reformulated as whether there exists $\varepsilon>0$ such that $e^{Dt}\cdot B(S,\varepsilon)$ does not intersect $e^{-Ct}\cdot T$ for all $t\geq 0$. We have thus reduced a question about a single linear dynamical system to a question about the interaction of two different (linear) dynamical systems governed by a rotation and a scaling matrix, respectively. This is what we mean by the decomposition method, and its study is one of the central objectives of this paper.
The sequence $e^{Dt}\cdot B(S,\varepsilon)$ is well-understood thanks to Kronecker’s classical theorem in Diophantine approximation. In particular, because $S$ is assumed to be bounded, for every $\varepsilon>0$ there exists a bounded set $H(\varepsilon)$ such that every open subset of $H(\varepsilon)$ intersects $e^{Dt}\cdot B(S,\varepsilon)$ for infinitely many (and unbounded) values of $t$. To analyse the sequence $(e^{-Ct}\cdot T)_{t\geq 0}$, on the other hand, we rely on o-minimality, which is a notion originating in model theory. We have that the first-order structure of real numbers equipped with arithmetic and real exponentiation is o-minimal, meaning that every subset of $\mathbb{R}^{d}$ that can be defined in first-order logic using arithmetic operations and real exponentiation has finitely many connected components (vdD1994bounded-analytic, ). We leverage this result to prove that $(e^{-Ct}\cdot T)_{t\geq 0}$ converges to a limit shape $L$ in a strong sense. The punchline of the argument is that $\langle M,S,T\rangle$ is a positive instance of the Robust Safety Problem if and only if (i) $L$ is disjoint from $H(\varepsilon)$ for all sufficiently small $\varepsilon$, and (ii) $e^{Dt}\cdot B(S,\varepsilon)$ does not intersect $e^{-Ct}\cdot T$ in the “transient phase”, i.e. for all values of $t$ up to an effectively computable time bound $N$. The value of $N$ is such that $e^{-Ct}\cdot T$ is sufficiently close to $L$ for all $t\geq N$.
The second problem we consider is the Pseudo-Orbit Safety Problem. In this setting, a perturbation is continuously applied to the trajectory at all times. Let $M\in\mathbb{R}^{d\times d}$ and $u\colon\mathbb{R}_{\geq 0}\to U$ be piecewise continuous for some $X\subseteq\mathbb{R}^{d}$. We say that $u$ is an input signal over $U$. In the presence of $u$, the system is evolves according to the dynamics $x^{\prime}(t)=Mx(t)+u(t)$. Hence the trajectory of $s$ in this case is given by
[…]
for all $t\geq 0$. We denote the right-hand side of the equality above by $\mathcal{T}(M,s,u,t)$. For $\varepsilon>0$, an $\varepsilon$-pseudo-orbit of $s$ (under $M$) is a trajectory of $s$ in the presence of an input signal over $B(\varepsilon)$, i.e. the open $\ell_{2}$-ball around the origin of radius $\varepsilon$. Pseudo-orbits are a fundamental notion in dynamical systems theory, going back to the works of Anosov (anosov, ), Bowen (bowen1975equilibrium, ), and Conley (conley1978isolated, ). For $M\in\mathbb{R}^{d\times d}$, $U\subseteq\mathbb{R}^{d}$, and $t\geq 0$, let
[…]
$\mathcal{R}(M,S,U,t)$$\coloneqq\{\mathcal{T}(M,s,u,t)\colon s\in S,u\textrm{ an input signal over $U$}\}=e^{Mt}\cdot S+\mathcal{P}(M,U,t).$
Then $\mathcal{R}(M,S,U,t)$ is the set of all points that are reachable from a point in $S$ at time exactly $t$ in the presence of an input signal from $U$. The Pseudo-Orbit Safety Problem for continuous-time LDS is to decide, given $M\in\mathbb{R}^{d\times d}$ with real algebraic entries and semialgebraic $S,T\subseteq\mathbb{R}^{d}$, whether there exists $\varepsilon>0$ such that $\mathcal{R}(M,S,B(\varepsilon),t)$ does not intersect $T$ for all $t\geq 0$. That is, whether there exists $\varepsilon>0$ such that for all $s\in S$, every $\varepsilon$-pseudo-orbit of $s$ avoids $T$.
In the discrete-time setting, an input signal $u$ has the type $\mathbb{N}\to\mathbb{R}^{d}$, and the system in the presence of $u$ evolves according to the relation $x(n+1)=Mx(n)+u(n)$. By induction it can be shown that the configuration of the system at time $n$ starting at $s$ is equal to $M^{n}s+\sum_{i=0}^{n-1}M^{i}u(n-1-i)$, which we denote by $\mathcal{T}_{d}(M,s,u,n)$. Let
$\mathcal{R}_{d}(M,S,U,n)\coloneqq\{\mathcal{T}_{d}(M,s,u,n)\colon s\in S,\textrm{ $u$ an input signal over $U$}\}=M^{n}\cdot S+\sum_{i=0}^{n-1}M^{i}\cdot U$
for $n\in\mathbb{N}$, which is the set of points reachable in the discrete-time LDS defined by $M$ from a point in $S$ at time $n$ in the presence of a control signal over $U$. The Pseudo-Orbit Safety Problem for discrete-time LDS is to decide, given $M,S,T$ as before, whether there exists $\varepsilon>0$ such that $\mathcal{R}_{d}(M,S,B(\varepsilon),n)$ does not intersect $T$ for all $t\geq 0$. When studying the Pseudo-Orbit Safety Problem, we will still use the decomposition method, but will apply o-minimality in a more intricate way. Our results are as follows.
Theorem 3.
Assuming Schanuel’s conjecture, the Pseudo-Orbit Safety Problem for continuous-time LDS is decidable for bounded $S$ and diagonalisable $M$. Moreover, given a positive instance $\langle M,S,T\rangle$ with bounded $M$ and diagonalisable $S$, we can effectively compute $\tilde{\varepsilon}\in\mathbb{Q}_{>0}$ such that $\mathcal{T}(M,S,B(\varepsilon),t)\notin T$ for all $t\geq 0$.
Theorem 4.
The Pseudo-Orbit Safety Problem for continuous-time LDS is decidable for bounded $S$ and diagonalisable and invertible $M$. Moreover, give a positive instance $\langle M,S,T\rangle$, we can compute $\tilde{\varepsilon}\in\mathbb{Q}_{>0}$ such that $\mathcal{T}_{d}(M,S,B(\varepsilon),n)\notin T$ for all $n\in\mathbb{N}$.
Our solution to the Pseudo-Orbit Safety Problem can also be applied to a third variant of the Safety Problem, where perturbations are applied both at the beginning and continuously at all times. We discuss this in Section 8.
For each of our two problems, we will give a single solution that is applicable to both continuous-time and discrete-time systems. The main idea is that, provided that $M$ is invertible, by taking matrix logarithms we can compute $\widetilde{M}$ such that $M^{n}=e^{\widetilde{M}n}$ for all $n\in\mathbb{N}$. That is, we can embed the discrete trajectories arising from $M$ into continuous ones generated by $\widetilde{M}$. The invertibility assumption is necessary, as $e^{A}$ is invertible for every $A$. Usually, in the study of decision problems of LDS a single approach is not sufficient for both discrete-time and continuous-time systems. For example, the techniques used for proving decidability of the Safety Problem with singleton $S$ and hyperplane targets in low dimensions differ substantially between the discrete and continuous settings (mignotte-shorey-tijdeman-skolem, ; chonev2023zeros, ).
The Safety Problem as well as our two problems are, intuitively, about the topological separation between the trajectories starting at $S$ and the set $T$. We will discuss a hierarchy of separability notions in Section 8. For now, we mention that there is another notion of separability based on invariants. An invariant of a (linear) dynamical system is a set $I$ of points whose trajectories always remain in $I$. Given $M,S,T$ as above, if we can construct an invariant $I$ of the LDS defined by $M$ such that $S\subseteq I$ and $I\cap T=\emptyset$, then we have a certificate of safety: all trajectories starting at a point in $S$ do not leave $I$ and hence never reach $T$. Such $I$ is said to separate $S$ from $T$. The Invariant Problem for a class $\mathcal{I}$ of sets is to decide, given $M,S,T$ as above, whether there exists an invariant $I\in\mathcal{I}$ that separates $S$ from $T$. When $\mathcal{I}$ is the set of all algebraic sets (in $\mathbb{R}^{d}$ for some $d$), then for every $S$ and $M$ there exists a unique smallest invariant enclosing all trajectories starting in $S$. For discrete-time LDS and singleton $S$, this strongest invariant is effectively computable (hrushovski2023strongest, ). The works (almagor2022discrete-invariants, ; almagor2020cont-invariants, ), on the other hand, show that, for both discrete-time and continuous-time LDS, given $M$ with real algebraic entries, a singleton $S$, and a semialgebraic set $T$, it is decidable whether there exists an invariant $I$ separating $S$ form $T$ that can be defined using arithmetic and exponentiation. We will discuss the connection between such invariants and the Robust Safety Problem in Section 6.2.
Structure of the paper
We will first make the decomposition method explicit, and study properties of rotation and scaling matrices separately. Section 3 establishes the technical results necessary for Section 5.1, and Section 4 doe the same for Section 5.2. We will study our two decision problems in Sections 6 and 7. In Section 8 we will discuss our new results and link them to what was previously known about linear dynamical systems.
1.1. Related work
In the previous joint work (dcosta2022pseudo, ) with D’Costa et al. we showed decidability of the Pseudo-Orbit Safety Problem for both discrete-time and continuos-time systems, under the restrictions that $M$ be diagonalisable and $S$ be a singleton. Our analysis was based on an ad hoc use of o-minimality theory, and we gave two separate proofs for the two classes of LDS. In this work, we give a single powerful framework that additionally allows us to generalise from singleton to bounded $S$. This unification comes at the cost of the additional assumption of invertibility (recall that we can embed only invertible discrete-time LDS into a continuous-time LDS as described above) in Theorem 4; See (karimov_2023, , Chap. 9) for how to handle non-invertible but diagonalisable $M$ using specialised arguments. When proving Theorem 2, on the other hand, we are able to reduce the general case to the invertible one, after which we perform the aforementioned embedding.
In (akshay2024robustness, ), among other “inexact” decision problems (with respect to the initial configuration of the system) for linear dynamical systems, Akshay et al. consider the Robust Safety Problem for discrete-time systems under the restrictions that $S$ be a singleton and $T$ be a hyperplane or a halfspace. They show decidability in both cases. In another joint work (dcosta2021pseudo, ) with D’Costa et al. we show decidability of the Pseudo-Orbit Safety Problem for discrete-time LDS and hyperplane or halfspace targets, without the diagonalisability restriction. Both of these results rely crucially on the fact that a hyperplane (respectively, a halfspace) can be defined by a single linear equation or inequality. The techniques of these two papers are direct and based on analysis of exponential polynomials using the classical theory of linear recurrence sequences; We do not discuss them further in this paper.
As mentioned earlier, decidability of the Safety Problem $S$ has been proven in some cases. In (almagor2021first, ), Almagor et al. show decidability for $M\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ with $d\leq 3$ and semialgebraic $S,T$. Recently, Vahanwala (vahanwala2023robust, ) has shown decidability in low dimensions assuming $T$ is a halfspace or a hyperplane and $S$ is definable using certain quadratic forms.
Some connections between o-minimality and dynamical (or, more generally, cyber-physical) systems have already been established. In (lafferriere2000minimal, ), Lafferriere et al. show that o-minimal hybrid systems always admit a finite bisimulation. In (miller2011expansions, ), Miller gives a dichotomy result concerning tameness of expansions of certain o-minimal structures with trajectories of linear dynamical systems. In this work, we use o-minimality in conjunction with the decomposition method. This framework can also be used to give geometric proofs of various known results about linear dynamical systems, an example of which can be found in Section 8.
2. Mathematical background
2.1. Notation and conventions
[…]
$B(x,\varepsilon)=\{y\in\mathbb{R}^{d}\colon\|x-y\|<\varepsilon\},$
We abbreviate $B(\mathbf{0},\varepsilon)$ to $B(\varepsilon)$. For $X\subseteq\mathbb{R}^{d}$ and $\varepsilon\geq 0$, we define
$B(X,\varepsilon)=\begin{cases}\{x+y\colon x\in X,y\in B(\varepsilon)\}\quad\textrm{ if }\varepsilon>0\\ X\qquad\qquad\qquad\qquad\qquad\,\textrm{ if }\varepsilon=0.\end{cases}$
For $x\in\mathbb{R}^{d}$ and $Y\subseteq\mathbb{R}^{d}$, we define $d(x,Y)=\inf\{\|y-x\|\colon y\in Y\}$.
For sets of vectors $X,Y$, we write $X+Y$ for $\{x+y\colon x\in X,y\in Y\}$. For a matrix $M$ and a set $X$ of vector, we write $M\cdot X$ to mean $\{Mx\colon x\in X\}$. Finally, for a set $\mathcal{M}$ of matrices and a set $X$ of vectors we define $\mathcal{M}\cdot X=\{Mx\colon M\in\mathcal{M},x\in X\}$.
We will work with sequences that are indexed by a subset of $\mathbb{N}$ or $\mathbb{R}_{\geq 0}$. When the indexing variable is $n$, for example $(X(n))_{n\in\mathbb{N}}$, the former is the case. When the variable is $t$, as in $(X(t))_{t\geq N}$ where $N\in\mathbb{R}_{\geq 0}$, the latter is the case. We say that a sequence of sets avoids a set $T$ if the intersection of the union of all terms of the sequence with $T$ is empty.
We write $\operatorname{Cl}(X)$ for the closure of $X$ in some set that will be either explicitly stated or clear from the context. We will only be working with Euclidean topology and induced subset topologies. When saying that an object is effectively computable, we mean that a representation of the object is effectively computable, in a representation scheme that will be clear from the context. For $X_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$, we write $\operatorname{diag}(X_{1},\ldots,X_{d})$ for the block-diagonal matrix in $\mathbb{R}^{d\times d}$, where $d=d_{1}+\ldots+d_{k}$, constructed from $X_{1},\ldots,X_{d}$ respecting the order.
2.2. Logical theories
A structure $\mathbb{M}$ consists of a universe $U$, constants $c_{1},\ldots,c_{k}\in U$, predicates $P_{1},\ldots,P_{l}$ where each $P_{i}\subseteq U^{\mu(i)}$ for some $\mu(i)\geq 1$, and functions $f_{1},\ldots,f_{m}$ where each $f_{i}$ has the type $f_{i}\colon U^{\delta(i)}\to U$ for some $\delta(i)\geq 1$. By the language of the structure $\mathbb{M}$, written $\mathcal{L}_{\mathbb{M}}$, we mean the set of all well-formed first-order formulas constructed from symbols denoting the constants $c_{1},\ldots,c_{k}$, predicates $P_{1},\ldots,P_{l}$, and functions $f_{1},\ldots,f_{m}$, as well as the symbols ,,,,¬,$\forall$,$\exists$,$\land$,$\lor$,$\lnot$,=∀ , ∃ , ∧ , ∨ , ¬ , =. A theory is simply a set of sentences, i.e. formulas without free variables. The theory of the structure $\mathbb{M}$, written $\operatorname{Th}(\mathbb{M})$, is the set of all sentences in the language of $\mathbb{M}$ that are true in $\mathbb{M}$. A theory $\mathcal{T}$
  • is decidable if there exists an algorithm that takes a sentence $\varphi$ and decides whether $\varphi\in\mathcal{T}$, and
  • admits quantifier elimination if for every formula $\varphi$ with free variables $x_{1},\ldots,x_{n}$, there exists a quantifier-free formula $\psi$ with the same free variables such that the formula
    $\forall x_{1},\ldots,x_{n}\colon(\varphi(x_{1},\ldots,x_{n})\Leftrightarrow\psi(x_{1},\ldots,x_{n}))$
    belongs to $\mathcal{T}$.
We will be working, for the most part, with the following structures and their theories.
  • (a)
    Let $\mathbb{R}_{0}=\langle\mathbb{R};0,1,<,+,\cdot\rangle$, which is the ring of real numbers. We will denote the language of this structure by $\mathcal{L}_{or}$, called the language of the ordered rings. Observe that using the constants $0,1$ and the addition, we can obtain any constant $c\in\mathbb{N}$. Hence every atomic formula in $\mathcal{L}_{or}$ with $k$ free variables is equivalent to $p(x_{1},\ldots,x_{k})\sim 0$, where $p$ is a polynomial with integer coefficients and $\sim$ is either $>$ or the equality. By the Tarski-Seidenberg theorem, $\operatorname{Th}(\mathbb{R}_{0})$ admits quantifier elimination and is decidable.
  • (b)
    Let $\mathbb{R}_{\exp}=\langle\mathbb{R};0,1,<,+,\cdot,\exp\rangle$, the real numbers augmented with the exponentiation function. The theory of this structure is model-complete, meaning that quantifiers in any formula can be eliminated down to a single block of existential quantifiers, and decidable assuming Schanuel’s conjecture. In fact, decidability of $\operatorname{Th}(\mathbb{R}_{\exp})$ implies a weak form of Schanuel’s conjecture.
  • (c)
    Finally, let $\mathbb{R}_{\exp,\,\operatorname{bt}}$ denote the structure $\langle\mathbb{R};0,1,<,+,\cdot,\exp,\widetilde{\cos},\widetilde{\sin}\rangle$, which is $\mathbb{R}_{\exp}$ augmented with bounded trigonometric functions: For $x\in(-1,1)$, $\widetilde{\cos}(x)=\cos(x)$ and $\widetilde{\sin}(x)=\sin(x)$. For all other values of $x$, we have $\widetilde{\cos}(x)=\widetilde{\sin}(x)=0$. This structure subsumes both $\mathbb{R}_{0}$ and $\mathbb{R}_{\exp}$.
We say that a structure $\mathbb{S}$ expands $\mathbb{M}$ if $\mathbb{S}$ and $\mathbb{M}$ have the same universe [^5] This is not the usual definition; can also extend the universe. and every constant, function, and relation of $\mathbb{M}$ is also present in $\mathbb{S}$. We will only need structures expanding $\mathbb{R}_{0}$. MAYBE DEFINE O-MINIMALITY HERE? , meaning that every subset of $\mathbb{R}$ definable in this structure is a finite union of intervals.
A set $X\subseteq U^{d}$ is definable in a structure $\mathbb{M}$ if there exist $k\geq 0$, a formula $\varphi$ in the language of $\mathbb{M}$ with $d+k$ free variables, and $a_{1},\ldots,a_{k}\in U$ such that for all $x_{1},\ldots,x_{d}\in U$, $\varphi(x_{1},\ldots,x_{d},a_{1},\ldots,a_{k})$ is true if and only if $(x_{1},\ldots,x_{d})\in X$. We say that $X$ is definable in $\mathbb{M}$ without parameters if we can take $k=0$ above.
2.3. Semialgebraic sets and algebraic numbers
A set $X\subseteq\mathbb{R}^{d}$ is semialgebraic if it is definable in $\mathbb{R}_{0}$ without parameters. By quantifier elimination, every semialgebraic set can be written as a Boolean combination of polynomial equalities and inequalities with integer coefficients. We say that $Z\subseteq\mathbb{C}^{d}$ is semialgebraic if $\tilde{Z}=\{(x_{1},y_{1},\ldots,x_{d},y_{d})\colon(x_{1}+\bm{i}y_{1},\ldots,x_{d}+\bm{i}y_{d})\in Z\}$ is a semialgebraic subset of $\mathbb{R}^{2d}$. We represent $Z$ by a formula defining $\tilde{Z}$. A function $f\colon X\to Y$ with $X\subseteq\mathbb{R}^{d}$, $Y\subseteq\mathbb{R}^{m}$ is semialgebraic if its graph $\{(x,f(x))\colon x\in X\}\subseteq\mathbb{R}^{d+m}$ is semialgebraic.
A number $z\in\mathbb{C}$ is algebraic if there exists $p\in\mathbb{Z}[x]$ such that $p(\alpha)=0$. The set of all algebraic numbers is denoted by $\overline{\mathbb{Q}}$. We will be working with real algebraic numbers. We will represent $x\in\mathbb{R}\cap\overline{\mathbb{Q}}$ by a formula $\varphi$ defining $\{x\}$. In this representation, arithmetic on real algebraic numbers is straightforward, and first-order properties of a given number can be verified using the decision procedure for $\operatorname{Th}(\mathbb{R}_{0})$.
2.4. Real Jordan form
For $a,b\in\mathbb{R}$, define $\Lambda(a,b)=\begin{bmatrix}a&-b\\ b&a\end{bmatrix}$, and for $\theta\in\mathbb{R}$, let $R(\theta)=\Lambda(\sin(\theta),\cos(\theta))$. A Jordan block is a matrix
(1) $J=\begin{bmatrix}\Lambda&I\\ &\Lambda&\ddots\\ &&\ddots&I\\ &&&\Lambda\end{bmatrix}\in\mathbb{R}^{d\times d}$
where $I$ is an identity matrix, and either (i) $\Lambda,I\in\mathbb{R}^{1\times 1}$, or (ii) $\Lambda=\rho R(\theta)$ with $I\in\mathbb{R}^{2\times 2}$, $\rho\in\mathbb{R}_{\geq 0}$, and $\theta\in\mathbb{R}$. The multiplicity of $J$ is $k$ in case (ii) and $k/2$ in case (ii). A matrix $J$ is in real Jordan form if $J=\operatorname{diag}(J_{1},\ldots,J_{l})$ where each $J_{i}$ is a real Jordan block. Given $M\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$, we can compute $P,J\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ such that $J$ is in real Jordan form and $M=P^{-1}JP$; See, for example, (cai1994computing, ). For any $\tilde{J}$ obtained by permuting the blocks of $J$, we can compute $Q\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ such that $M=Q^{-1}\tilde{J}Q$.
2.5. Matrix exponential
For $M\in\mathbb{R}^{d\times d}$, let $e^{M}=\sum_{n=0}^{\infty}\frac{M^{n}}{n!}$. For all $M$, this series converges and $e^{M}$ is invertible. If $A,B$ commute, then $e^{A+B}=e^{A}e^{B}$. For $\Lambda=\Lambda(a,b)$ we have that $e^{\Lambda}=e^{a}R(b)$.
Suppose $M=P^{-1}JP$ with $P,J\in\mathbb{R}^{d\times d}$ and $J$ in real Jordan form. Further suppose $J=\operatorname{diag}(J_{1},\ldots,J_{l})$, where each $J_{i}=\operatorname{diag}(\Lambda_{i},\ldots,\Lambda_{i})$ is a Jordan block with multiplicity $\sigma_{i}$. Let $t\geq 0$. We have that $e^{Mt}=P^{-1}e^{Jt}P$, $e^{Jt}=\operatorname{diag}(e^{J_{1}t},\ldots,e^{J_{l}t})$, and for all $i$,
(2) $e^{J_{i}t}=\begin{bmatrix}e^{\Lambda_{i}t}&te^{\Lambda_{i}t}&\cdots&\frac{t^{\sigma}_{i}}{\sigma!}e^{\Lambda_{i}t}\\ &e^{\Lambda_{i}t}&\ddots&\vdots\\ &&\ddots&te^{\Lambda_{i}t}\\ &&&e^{\Lambda_{i}t}\end{bmatrix}.$
Now suppose $J$ is invertible. For $1\leq i\leq m$, define
$\log(\Lambda_{i})=\begin{cases}[\log(\rho_{i})],\qquad\quad\>\>\textrm{ if }\Lambda_{i}=[\rho_{i}]\\ \Lambda(\log(\rho_{i}),\theta),\quad\quad\textrm{ if }\Lambda_{i}=\rho_{i}R(\theta_{i}).\end{cases}$
We can then define the matrix logarithm of $J$ as $\log(J)=\operatorname{diag}(\log(J_{1}),\ldots,\log(J_{l}))$, where for all $i$,
$\log(J_{i})=\begin{bmatrix}\log(\Lambda_{i})&\Lambda_{i}^{-1}&\cdots&-\frac{(-\Lambda_{i}^{-1})^{\sigma_{k}}}{\sigma_{k}}\\ &\log(\Lambda_{i})&\ddots&\vdots\\ &&\ddots&\Lambda_{i}^{-1}\\ &&&\log(\Lambda_{i})\end{bmatrix}.$
With this definition, for all $n\in\mathbb{N}$, $M^{n}=P^{-1}e^{\log(J)n}P$; See (culver1966existence, ). We will need the following lemma about matrix exponentiation.
Lemma 1.
If $M$ is nilpotent, then for every $i,j$ there exists a polynomial $p$ with real algebraic coefficients such that $(e^{Mt})_{i,j}=p(t)$ for all $t\geq 0$.
Proof.
Write $M=P^{-1}JP$, where $J$ is in real Jordan form. By the nilpotence assumption, $J=\operatorname{diag}(J_{1},\ldots,J_{l})$ where $J_{i}$ has ones on the superdiagonal and zeros everywhere else, i.e. $\Lambda_{i}=[0]$ for all $i$ in the notation above. It remains to recall (2). ∎
3. Applications of Kronecker’s theorem
We now discuss Kronecker’s classical theorem in simultaneous Diophantine approximation, which will play a crucial role in our analysis of dynamical systems given by a rotation matrix.
Theorem 1 (Kronecker, see (bourbaki1966elements, , Chap. 7, Sec. 1.3, Prop. 1.7)).
Let $\lambda_{1},\ldots,\lambda_{l}$ and $x_{1},\ldots,x_{l}$ be real numbers such that for all $u_{1},\ldots,u_{l}\in\mathbb{Z}$,
$u_{1}\lambda_{1}+\cdots+u_{l}\lambda_{l}\in\mathbb{Z}\Rightarrow u_{1}x_{1}+\cdots+u_{l}x_{l}\in\mathbb{Z}.$
Then for every $\varepsilon>0$ there exists $n\in\mathbb{N}$ such that $|n\lambda_{i}-x_{i}|<\varepsilon$ for all $i$.
Corollary 2.
Let $\lambda_{1},\ldots,\lambda_{l},x_{1},\ldots,x_{l}$ be as above. For every $\varepsilon>0$, there exist infinitely many $n\in\mathbb{N}$ such that $|n\lambda_{i}-x_{i}|<\varepsilon$ for all $i$.
Proof.
For all $k\in\mathbb{N}$ and $u_{1},\ldots,u_{l}\in\mathbb{Z}$,
$u_{1}\lambda_{1}+\cdots+u_{l}\lambda_{l}\in\mathbb{Z}\Rightarrow u_{1}(x_{1}+k\lambda_{1})+\cdots+u_{l}(x_{l}+k\lambda_{l})\in\mathbb{Z}.$
By Theorem 1, for every $k$ there exist $n_{k}\in\mathbb{N}$ such that $|n_{k}\lambda_{i}-(x_{i}+k\lambda_{i})|<\varepsilon$ for all $i$. Hence $|(n_{k}+k)\lambda_{i}-x_{i}|<\varepsilon$ for all $k$ and $i$. It remains to observe that $n_{k}+k$ grows arbitrarily large as $k\to\infty$. ∎
Let $b_{1},\ldots,b_{l}\in\mathbb{R}$, $\beta=(e^{\bm{i}b_{1}},\ldots,e^{\bm{i}b_{l}})\in\mathbb{T}^{l}$ and $\beta^{t}=(e^{\bm{i}b_{1}t},\ldots,e^{\bm{i}b_{l}t})$ for $t\geq 0$. We will use Kronecker’s theorem to analyse the trajectory $(\beta^{t})_{t\geq 0}$ in $\mathbb{T}^{l}$. Define
$G(\beta)=\{(k_{1},\ldots,k_{l})\in\mathbb{Z}^{l}\colon e^{\bm{i}k_{1}b_{1}}\cdots e^{\bm{i}k_{l}b_{l}}=1\}$
and observe that $G(\beta)$ is an abelian group with a finite basis $V=\{v_{1},\ldots,v_{m}\}$ where $m\leq l$ and $v_{i}\in\mathbb{Z}^{l}$ for all $i$. Write $v_{i}=(v_{i,1},\ldots,v_{i,l})$ for $1\leq i\leq m$. Next, define
$\mathcal{X}(\beta)=\{z\in\mathbb{T}^{l}\colon G(\beta)\subseteq G(z)\}\subseteq\mathbb{T}^{l}.$
The set $\mathcal{X}(\beta)\subseteq\mathbb{T}^{l}$ is compact and semialgebraic, as it is the set of all $(z_{1},\ldots,z_{l})\in\mathbb{T}^{l}$ such that $z_{1}^{i,1}\cdots z_{1}^{i,l}=1$ for all $1\leq i\leq m$. Moreover, a representation of $\mathcal{X}(\beta)$ can be effectively computed given a basis of $G(\gamma)$. We next show that $(\beta^{t})_{t\geq 0}$ is uniformly recurrent in $\mathcal{X}(\beta)$.
Theorem 3.
Let $\alpha=(e^{\bm{i}a_{1}},\ldots,e^{\bm{i}a_{l}})\in\mathcal{X}(\beta)$. For every open subset $O$ of $\mathbb{T}^{l}$ containing $\alpha$ there exist infinitely many $n\in\mathbb{N}$ such that $\beta^{n}\in O$.
Proof.
For $x,y\in\mathbb{R}$ let $[\![x]\!]_{y}$ be the closest distance from $x$ to a multiple of $y$, and $[\![x]\!]=[\![x]\!]_{1}$. For all $n\in\mathbb{N}$,
$\|\beta^{n}-z\|_{\infty}=\max_{1\leq i\leq l}|e^{\bm{i}nb_{i}}-e^{\bm{i}a_{i}}|\leq\max_{1\leq i\leq l}\,[\![nb_{i}-a_{i}]\!]_{2\pi}=2\pi\cdot\max_{1\leq i\leq l}\,[\![nb_{i}/(2\pi)-a_{i}/(2\pi)]\!].$
We will apply Kronecker’s theorem to the right-hand side of the last equality. Let $u_{1},\ldots,u_{l}\in\mathbb{Z}$, and suppose
$\frac{u_{1}b_{1}}{2\pi}+\ldots+\frac{u_{l}b_{l}}{2\pi}\in\mathbb{Z}$
which is equivalent to $e^{\bm{i}u_{1}b_{1}}\cdots e^{\bm{i}u_{l}b_{l}}=1$. Since $\alpha\in X(\beta)$, $G(\beta)\subseteq G(\alpha)$ and hence $e^{\bm{i}u_{1}a_{1}}\cdots e^{\bm{i}u_{l}a_{l}}=1$. It follows that $u_{1}a_{1}+\ldots+u_{l}a_{l}\in\mathbb{Z}$. Applying Corollary 2, for every $\varepsilon>0$ there exist infinitely many $n\in\mathbb{N}$ such that $|nb_{i}/(2\pi)-a_{i}/(2\pi)|<\varepsilon$, which, for $\varepsilon\in(0,1/2)$, implies that $[\![nb_{i}/(2\pi)-a_{i}/(2\pi)]\!]<\varepsilon$. Hence for every $\varepsilon\in(0,1/2)$ there exist infinitely many $n\in\mathbb{N}$ such that $\|\beta^{n}-z\|_{\infty}<\varepsilon$. It remains to construct, for the given $O$, a value $\varepsilon>0$ such that $\{z\in\mathbb{T}^{l}\colon\|z-\alpha\|_{\infty}<\varepsilon\}\subseteq O$. ∎
Corollary 4.
The set $\mathcal{X}(\beta)$ is the topological closure (both in $\mathbb{T}^{l}$ and $\mathbb{C}^{l}$) of $(\beta^{t})_{t\geq 0}$.
Proof.
For every $t\geq 0$, $G(\beta)\subseteq G(\beta^{t})$ and hence $\beta^{t}\in\mathcal{X}(\beta)$. By Theorem 3, $\mathcal{X}(\beta)$ is the closure of $(\beta^{t})_{t\geq 0}$ in $\mathbb{T}^{l}$. It remains to observe that $\mathbb{T}^{l}$ is closed in $\mathbb{C}^{l}$. It remains to invoke Theorem 3. ∎
We conclude by giving two theorems that show that, under certain assumptions on $b_{1},\ldots,b_{l}$, a basis for $G(\beta)$ and hence a representation of $\mathcal{X}(\beta)$ can be effectively computed. These theorems respectively correspond to continuous-time and discrete-time LDS given by $M\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$.
Theorem 5.
Suppose $e^{\bm{i}b_{i}}\in\overline{\mathbb{Q}}$ for all $i$. Then a representation of $\mathcal{X}(\beta)$ can be effectively computed.
Proof.
In this case, $G(\beta)=\{(k_{1},\ldots,k_{l})\in\mathbb{Z}^{l}\colon\rho_{1}^{k_{1}}\cdots\rho_{l}^{k_{l}}=1\}$, where $\rho_{i}=e^{\bm{i}b_{i}}$ for all $i$. Masser (masser-mult-rel-bound, ) gives a bound $M$ that can be effectively computed from representations of $\rho_{1},\ldots,\rho_{l}$ such that $G(\beta)$ has a basis $\{v_{1},\ldots,v_{m}\}$ satisfying $\|v_{j}\|_{\infty}<M$ for all $i$. To compute a basis of $G(\beta)$, start with $V=\emptyset$ and construct increasingly larger linearly independent (over $\mathbb{Z}$) subsets of $G(\beta)$. At each step, consider fresh $v=(v_{1},\ldots,v_{l})\in\mathbb{Z}^{l}$ with $\|v\|_{\infty}<M$. For this $v$, check whether $\rho_{1}^{v_{1}}\cdots\rho_{l}^{v_{l}}=1$, that is, whether $v\in G(\beta)$. If this is the case and $V\cup\{v\}$ is linearly independent, then add $v$ to $V$. The algorithm terminates since a basis of $G(\beta)$ contains at most $l$ vectors. ∎
Theorem 6.
Suppose $b_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$. Then a representation of $\mathcal{X}(\beta)$ can be effectively computed.
Proof.
For all $\lambda\in\mathbb{R}\cap\overline{\mathbb{Q}}$, $e^{\bm{i}\lambda}=1$ if and only $\lambda=0$. Hence $G(\beta)=\{(k_{1},\ldots,k_{l})\in\mathbb{Z}^{l}\colon k_{1}b_{1}+\ldots+k_{l}b_{l}=0\}$. Let $\mathbb{K}=\mathbb{Q}(b_{1},\ldots,b_{l})$ and $D$ be the degree of the field extension $\mathbb{K}/\mathbb{Q}$. Using the Primitive Element Theorem (cohen2013course, , Chap. 4.2.2), we can compute $\theta\in\mathbb{K}$ such that $\varphi\colon\mathbb{Z}^{D}\to\mathbb{K}$ given by $\varphi(x_{0},\ldots,x_{D-1})=\sum_{j=0}^{D-1}x_{j}\theta^{j-1}$ is a bijection. By considering $\varphi^{-1}(b_{i})$ for $1\leq i\leq l$, we can re-write $k_{1}b_{1}+\ldots+k_{l}b_{l}=0$ as $\sum_{j=0}^{D-1}L_{j}(k_{1},\ldots,k_{l})\theta^{j}=0$, where each $L_{j}$ is a linear form with integer coefficients. Since $\sum_{j=0}^{D-1}x_{j}\theta^{j-1}=0$ if and only if $x_{j}=0$ for all $j$, the abelian group $G(\beta)$ is defined by the linear equations (with integer coefficients) $L_{j}(k_{1},\ldots,k_{l})=0$ for $j=1,\ldots,D$. We can therefore compute a basis for $G(\beta)$ using, for example, the Hermite normal form (cohen2013course, , Chap. 2.4). ∎
4. Real exponentiation and o-minimality
Let $g_{1},\ldots,g_{k}\colon\mathbb{R}\to\mathbb{R}$ be real analytic over $(-1,1)$, and define, for $1\leq i\leq k$,
$f_{i}(x)=\begin{cases}g_{i}(x)\quad\>\>\textrm{ if $0<x<1$},\\ 0\quad\qquad\textrm{ otherwise}.\end{cases}$
By (vdD1994bounded-analytic, ), the structure $\langle\mathbb{R};0,1,<,+,\cdot,\exp,f_{1},\ldots,f_{k}\rangle$ is o-minimal. In particular, the structures $\mathbb{R}_{0}$, $\mathbb{R}_{\exp}$, and $\mathbb{R}_{\exp,\,\operatorname{bt}}$ are all o-minimal. As mentioned in the introduction, we will make certain (effective) convergence arguments in o-minimal structures that will be critical to our analysis of dynamical systems given by a scaling matrix.
We say that a sequence $(Z_{t})_{t\geq 0}$ of subsets of $\mathbb{R}^{d}$ is definable in a structure $\mathbb{S}$ expanding $\mathbb{R}_{0}$ if there exist $k\geq 0$, $a_{1},\ldots,a_{k}\in\mathbb{R}$, and $\varphi\in\mathcal{L}_{\mathbb{S}}$ with $k+d+1$ free variables such that for all $t\geq 0$ and $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$, $x\in Z_{t}$ if and only if $\mathbb{S}\models\varphi(x_{1},\ldots,x_{d},t,a_{1},\ldots,a_{k})$. The following definition is an adaptation of Kuratowski convergence to sequences over $\mathbb{R}_{\geq 0}$, and captures the kind of set-to-set convergence with which we will work.
Definition 1.
Let $(Z_{t})_{t\geq 0}$ be a sequence of subsets of $\mathbb{R}^{d}$, and $L=\{x\in\mathbb{R}^{d}\colon\liminf_{t\to\infty}d(x,Z_{t})=0\}$. The sequence $(Z_{t})_{t\geq 0}$ converges if $L=\{x\in\mathbb{R}^{d}\colon\lim_{t\to\infty}d(x,Z_{t})=0\}$, in which case $L$ is is called the limit shape of $(Z_{t})_{t\geq 0}$.
We have the following.
Lemma 2.
Suppose $L$ is the limit shape of $(Z_{t})_{t\geq 0}$. Then
  • (a)
    $L$ is closed,
  • (b)
    $L=\emptyset$ if and only if for every compact $X$, there exists $N$ such that $Z_{t}\cap X=\emptyset$ for all $t\geq N$, and
  • (c)
    for every compact $X\subseteq\mathbb{R}^{d}$ and $\varepsilon>0$, there exists $N\geq 0$ such that for all $t\geq N$,
    $Z_{t}\cap X\subseteq B(L\cap X,\varepsilon).$
Proof.
Statement (a) is immediate from the definition, and (b) follows from the Bolzano-Weierstraß theorem. To prove (c), fix compact $X$ and $\varepsilon>0$. The set $Y\coloneqq X\setminus B(L\cap X)$ is compact. For every $y\in Y$, since $Y\notin L$, there exist $T_{y}\geq 0$ and an open subset $B_{y}\ni y$ of $Y$ such that $Z_{t}\cap B(y)=\emptyset$ for all $t\geq T_{y}$. Hence $\mathcal{C}=\{B_{y}\colon y\in Y\}$ is an open cover of $Y$. Let $\mathcal{F}$ be a finite sub-cover. We can then take $T=\max\{T_{y}\colon y\in\mathcal{F}\}$. ∎
Note that by (c) above, our definition of convergence is the same as convergence with respect to the Hausdorff metric. Next, recall that a function definable in an o-minimal structure is ultimately monotonic (vdD-geometric-categories, , Sec. 4.1). The following is an immediate consequence, but we give a little more detail to illustrate the role of o-minimality.
Theorem 3.
Every $(Z_{t})_{t\geq 0}$ definable in an o-minimal expansion $\mathbb{S}$ of $\mathbb{R}_{0}$ converges.
Proof.
Let $a_{1},\ldots,a_{k}\in\mathbb{R}$ and $\varphi\in\mathcal{L}_{\mathbb{S}}$ with $m+k+1$ free variables be such that $\varphi(a_{1},\ldots,a_{m},t,x_{1},\ldots,x_{k})$ defines $(Z_{t})_{t\geq 0}$ in $\mathbb{S}$. Consider $x\in\mathbb{R}^{k}$ with $\lim_{t\to\infty}d(x,Z_{t})=0$. Let $\Delta>0$ and $T=\{t\geq 0\colon d(x,Z_{t})<\Delta\}$, which is definable in $\mathbb{S}$. By o-minimality, $T$ is a finite union of interval subsets of $\mathbb{R}_{\geq 0}$. Since $\liminf_{t\to\infty}d(x,Z_{t})=0$, $T$ is unbounded. Therefore, $T$ must enclose an interval of the form $[N,\infty)$. That is, $d(x,Z_{t})<\Delta$ for all sufficiently large values of $t$. It follows that $\lim_{t\geq 0}d(x,Z_{t})=0$. ∎
In the cases we will encounter, the limit shape $L$ will be a semialgebraic set whose representation can be computed effectively. To show this, we will need the following lemma. Let $\mathbb{E}=(\mathbb{R}_{>0}\cap\overline{\mathbb{Q}})\cup\{e^{x}\colon x\in\mathbb{R}_{>0}\cap\overline{\mathbb{Q}}\}$. This set in particular contains the eigenvalues of both discrete-time and continuous-time dynamical systems governed by a scaling matrix $M\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$.
Lemma 4.
Let $\varphi\in\mathcal{L}_{or}$ with $k+m+1$ free variables, and $\rho_{1},\ldots,\rho_{m}\in\mathbb{E}$. Define
$A=\{(x_{1},\ldots,x_{k})\in\mathbb{R}^{k}\mid\exists N.\,\forall t\in[N,\infty)\colon\varphi(x_{1},\ldots,x_{k},t,\rho_{1}^{t},\ldots,\rho_{m}^{t})\}$
and
$B=\{(x_{1},\ldots,x_{k})\in\mathbb{R}^{k}\mid\exists N.\,\forall t\in[N,\infty)\colon\lnot\varphi(x_{1},\ldots,x_{k},t,\rho_{1}^{t},\ldots,\rho_{m}^{t})\}.$
We have the following.
  1. (a)
    $A\cup B=\mathbb{R}^{k}$.
  2. (b)
    Both $A$ and $B$ are semialgebraic sets whose representations can be computed effectively.
  3. (c)
    Given $(x_{1},\ldots,x_{k})\in B\cap\mathbb{Q}^{k}$, we can effectively compute $N\in\mathbb{N}$ such that $\varphi(x_{1},\ldots,x_{k},t,\rho_{1}^{t},\ldots,\rho_{m}^{t})$ does not hold for all $t\geq N$.
Proof.
Let $(x_{1},\ldots,x_{k})\in\mathbb{R}^{k}$, and define $T=\{t\geq 0\colon\varphi(x_{1},\ldots,x_{k},t,\rho_{1}^{t},\ldots,\rho_{m}^{t})\}$. By o-minimality of $\mathbb{R}_{\exp}$, either $T$ is bounded, or it encloses an interval of the form $[N,\infty)$. That is, $(x_{1},\ldots,x_{k})$ belongs to either $A$ or $B$. This proves (a).
Next, using quantifier elimination in $\operatorname{Th}(\mathbb{R}_{0})$, compute a formula
(3) $\bigvee_{i\in I}\bigwedge_{j\in J}p_{i,j}(x_{1},\ldots,x_{k},y_{0},\ldots,y_{m})\mathrel{\Delta_{i,j}}0$
in $\mathcal{L}_{or}$ equivalent to $\varphi(x_{1},\ldots,x_{k},y_{0},\ldots,y_{m})$, where each $p_{i,j}$ is a polynomial with rational coefficients and $\Delta_{i,j}\in\{>,=\}$. Fix $i\in I$ and $j\in J$. Observe that for every $r_{1},\ldots,r_{m}\in\mathbb{N}$, $\rho_{1}^{r_{1}}\cdots\rho_{m}^{r_{m}}$ is of the form $\lambda e^{\mu}$ where $\lambda,\mu\in\mathbb{R}_{>0}\cap\overline{\mathbb{Q}}$. By the Gelfond-Schneider theorem (waldschmidt2013diophantine, , Chap. 1.4), for all $\lambda_{1},\mu_{1},\lambda_{2},\mu_{2}\in\overline{\mathbb{Q}}$, $\lambda_{1}e^{\mu_{1}}=\lambda_{2}e^{\mu_{2}}$ if and only if $\lambda_{1}=\lambda_{2}$ and $\mu_{1}=\mu_{2}$. Therefore, we can effectively write
$p_{i,j}(x_{1},\ldots,x_{k},t,\rho_{1}^{t},\ldots,\rho_{m}^{t})=\sum_{l=1}^{d}h_{l}(x_{1},\ldots,x_{k})q_{l}(t)R_{l}^{t}$
where $R_{1}>\ldots>R_{d}>0$, each $R_{i}$ is of the form $\lambda_{i}e^{\mu_{i}}$ for $\lambda_{i},\mu_{i}\in\mathbb{R}_{>0}\cap\overline{\mathbb{Q}}$, and $h_{l},q_{l}$ are non-zero polynomials. Without loss of generality we can further assume that $q_{l}(t)>0$ for all sufficiently large $t$. Let
$A_{i,j}=\{(x_{1},\ldots,x_{k})\in\mathbb{R}^{k}\colon p_{i,j}(x_{1},\ldots,x_{k},t,r_{1}^{t},\ldots,r_{m}^{t})\mathrel{\Delta_{i,j}}0\textrm{ for all sufficiently large $t$}\}.$
Observe that whether $(x_{1},\ldots,x_{k})\in A_{i,j}$ only depends on the sign of $h_{l}(x_{1},\ldots,x_{k})$ for $1\leq l\leq d$. In particular, if $\Delta_{i,j}$ is $=$, then $A_{i,j}$ is defined by $\bigwedge_{l=1}^{d}h_{l}(x_{1},\ldots,x_{k})=0$. If $\Delta_{i,j}$ is $>$, then $A_{i,j}$ defined by
$\bigvee_{l=1}^{d}\biggl{(}h_{l}(x_{1},\ldots,x_{k})>0\land\bigwedge_{s=1}^{l-1}h_{s}(x_{1},\ldots,x_{k})=0\biggr{)}.$
Hence $A_{i,j}$ is semialgebraic with an effectively computable representation. It remains to observe that $A=\bigvee_{i\in I}\bigwedge_{j\in J}A_{i,j}$.
To prove (c), let $x=(x_{1},\ldots,x_{k})\in B\cap\mathbb{Q}^{k}$. Let $\varphi\in\mathcal{L}_{or}$ be the formula obtained by substituting the values of $x_{1},\ldots,x_{k}$ into (3), which will be of the form
$\psi(y_{0},\ldots,y_{m})=\bigvee_{i\in I}\bigwedge_{j\in J}v_{i,j}(y_{0},\ldots,y_{m})\mathrel{\Delta_{i,j}}0$
for polynomials $v_{i,j}$ with rational coefficients. We have to construct $N$ such that for all $t\geq N$, $\psi(t,r_{1}^{t},\ldots,r_{m}^{t})$ does not hold. To do this, it suffices to construct, for each $i\in I$ and $j\in J$, an integer $N_{i,j}$ such that either $v_{i,j}(t,r_{1}^{t},\ldots,r_{m}^{t})\mathrel{\Delta_{i,j}}0$ holds for all $t\geq N_{i,j}$, or it does not hold for all $t\geq N_{i,j}$. We can then take $N=\max_{i,j}N_{i,j}$.
Fix $i\in I$, $j\in J$, and write $v_{i,j}(t,r_{1}^{t},\ldots,r_{m}^{t})=\sum_{l=1}^{d}q_{l}(t)R_{l}^{t}$ where $R_{1}>\ldots>R_{d}>0$, each $R_{i}$ is of the form $\lambda_{i}e^{\mu_{i}}$ for $\lambda_{i},\mu_{i}\in\mathbb{R}_{>0}\cap\overline{\mathbb{Q}}$, and $q_{l}$ is a non-zero polynomial with rational coefficients. Compute rationals $M_{1},M_{2},c>0$ and $N_{i,j}\in\mathbb{N}$ such that $R_{1}>M_{1}>M_{2}>R_{2}$, and for all $t\geq N_{i,j}$, $|q_{1}(t)|>c$ and $cM_{1}^{t}>(d-1)q_{l}(t)M_{2}^{t}$. Then the sign of $v_{i,j}(t,r_{1}^{t},\ldots,r_{m}^{t})$ is stable over $[N_{i,j},\infty)$. ∎
The following is our main effective convergence result.
Lemma 5.
Let $(Z_{t})_{t\geq 0}$ be a sequence of subsets of $\mathbb{R}^{d}$. Suppose there exist $\varphi\in\mathcal{L}_{or}$ and $\rho_{1},\ldots,\rho_{d}\in\mathbb{E}$ such that for all $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ and $t\geq 0$,
(4) $x\in Z_{t}\Leftrightarrow\varphi(x_{1},\ldots,x_{d},t,\rho_{1}^{t},\ldots,\rho_{d}^{t}).$
Then the limit shape $L$ of $(Z_{t})_{\geq 0}$ is semialgebraic and can be effectively computed from $\varphi,\rho_{1},\ldots,\rho_{d}$.
Proof.
By (4) and the fact that each $\rho_{i}$ is definable in $\mathbb{R}_{\exp}$, the sequence $(e^{Ct}\cdot T)_{t\geq 0}$ is definable in $\mathbb{R}_{\exp}$. By Theorem 3, the sequence converges to a limit shape $L$. By Lemma 4,
$X=\{(x_{1},\ldots,x_{d},\varepsilon)\mid\varepsilon>0\textrm{ and }\exists N.\>\forall t\in[N,\infty)\colon d((x_{1},\ldots,x_{d}),e^{Ct}\cdot T)<\varepsilon\}$
is semialgebraic and effectively computable. It remains to observe that the limit shape is
$L=\{(x_{1},\ldots,x_{d})\mid\forall\varepsilon>0\colon(x_{1},\ldots,x_{d},\varepsilon)\in X\}$
which is semialgebraic and can be effectively computed from $X$. ∎
5. The decomposition method
Consider $M\in\mathbb{R}^{d\times d}$ with eigenvalues $\lambda_{1},\ldots,\lambda_{d}$. We say that $M$ s a scaling matrix if $\lambda_{1},\ldots,\lambda_{d}\in\mathbb{R}_{\geq 0}$, and $M$ is a rotation matrix if it is diagonalisable and $\lambda_{1},\ldots,\lambda_{d}\in\mathbb{T}$, and $M$. Observe that $e^{Mt}$ is a rotation matrix for all $t\geq 0$ if and only if $M$ is diagonalisable and $\lambda_{j}\in\bm{i}\>\!\mathbb{R}$ for all $j$. Alternatively, $e^{Mt}$ is a rotation matrix for all $t\geq 0$ if and only if $M$ can be written as $P^{-1}JP$ where $J=\operatorname{diag}(\Lambda_{1},\ldots,\Lambda_{l},\Pi)$ is in real Jordan form, $\Lambda_{i}=\Lambda(0,b_{i})$ for all $i$, and $\Pi$ is a (possibly empty) zero matrix. On the other hand, $e^{Mt}$ is a scaling matrix for all $t\geq 0$ if and only if $M$ is a scaling matrix.
Definition 1.
A decomposition of $M\in\mathbb{R}^{d\times d}$ is a pair $(C,D)$ of matrices with $C,D\in\mathbb{R}^{d\times d}$ such that $M=C+D$, all eigenvalues of $C$ are real, $D$ is diagonalisable with only purely imaginary eigenvalues, and $C,D$ commute.
We will later show that a decomposition always exists, and can be effectively computed if the entries of $M$ are real algebraic. For the moment, let $(C,D)$ be a decomposition of $M$. We have that for all $t\geq 0$, $e^{Ct}$ is a scaling matrix, $e^{Dt}$ is a rotation matrix, and $e^{Mt}=e^{Ct}e^{Dt}$. The last property follows from the fact that $e^{A+B}=e^{A}e^{B}$ for commuting $A,B$. Observe that for all $t\geq 0$, $e^{-Ct}$ is a scaling matrix that is the inverse of $e^{Ct}$. By the decomposition method we refer to studying whether $e^{Mt}\cdot X$ intersects $T$ (where $X,T\subseteq\mathbb{R}^{d}$) by considering whether $e^{Dt}\cdot X$ intersects $e^{-Ct}\cdot T$. Using this method we can, for example, recast the Robust Safety Problem as the problem of deciding whether there exists $\varepsilon>0$ such that $e^{Dt}\cdot B(S,\varepsilon)$ does not intersect $e^{-Ct}\cdot T$ for all $t\geq 0$. The advantage of the latter view, as we will see shortly, is that the sequences of sets $(e^{Dt}\cdot X)_{t\geq 0}$ and $(e^{-Ct}\cdot T)_{t\geq 0}$ separately follow two different kinds of dynamics that have many desirable properties.
5.1. Dynamics of rotation matrices
Consider diagonalisable $D\in\mathbb{R}^{d\times d}$ with only purely imaginary eigenvalues. We next study the sequence $(e^{Dt})_{t\geq 0}$ of rotation matrices. Let $\mathcal{D}\subseteq\mathbb{R}^{d\times d}$ be the topological closure of $\{e^{Dt}\colon t\geq 0\}$.
Lemma 2.
We have the following.
  • (a)
    The set $\mathcal{D}$ is compact and definable in $\mathbb{R}_{0}$.
  • (b)
    For every $A\in\mathcal{D}$, $\det(A)=1$. Moreover, there exists $b>0$ such that $\|A\|,\|A^{-1}\|\leq b$ for all $A\in\mathcal{D}$.
  • (c)
    For every non-empty open subset $O$ of $\mathcal{D}$, there exist infinitely many $n\in\mathbb{N}$ such that $e^{Dn}\in O$.
  • (d)
    Suppose $D=P^{-1}JP$, where $P\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$,
    $J=\operatorname{diag}(\Lambda(0,b_{1}),\ldots,\Lambda(0,b_{m}),\mathbf{0})$
    and either (i) $b_{1},\ldots,b_{m}\in\mathbb{R}\cap\overline{\mathbb{Q}}$, or (ii) $b_{1},\ldots,b_{m}\in\{\operatorname{Log}(z)/\bm{i}\colon z\in\mathbb{T}\cap\overline{\mathbb{Q}}\}$. Then $\mathcal{D}$ is semialgebraic and effectively computable.
Proof.
We first prove (a). Write $D=P^{-1}JP$, where $J=\operatorname{diag}(\Lambda(0,b_{1}),\ldots,\Lambda(0,b_{m}),\mathbf{0})$; This is always possible due to the assumptions on the spectrum of $D$. Let
$U=\{\operatorname{diag}(R(\theta_{1}),\ldots,R(\theta_{m}),I)\colon\theta_{1},\ldots,\theta_{m}\in\mathbb{R}\}\subseteq\mathbb{R}^{d\times d}$
and $f(e^{\bm{i}\theta_{1}},\ldots,e^{\bm{i}\theta_{m}})=\operatorname{diag}(R(\theta_{1}),\ldots,R(\theta_{m}),I)$. We have that $f$ is a semialgebraic homeomorphism (see Section 2.3) between $\mathbb{T}^{m}$ and $U$, and $f(e^{\bm{i}b_{1}t},\ldots,e^{\bm{i}b_{m}t})=e^{Jt}$ for all $t\geq 0$. Let $\mathcal{J}$ be the closure of $(e^{Jt})_{t\geq 0}$. Since $U\supseteq\mathcal{J}$ and $U$ is closed in $\mathbb{R}^{d\times d}$, $\mathcal{J}=f(\mathcal{X})$ where $\mathcal{X}$ is the closure of $\{(e^{\bm{i}b_{1}t},\ldots,e^{\bm{i}b_{m}t})\colon t\geq 0\}$ in $\mathbb{T}^{l}$. Observe that $\mathcal{D}=P^{-1}\mathcal{J}P$. Since $\mathcal{X}$ is compact, $\mathcal{J}$ and $\mathcal{D}$ are also compact. As discussed in Section 3, $\mathcal{X}\subseteq\mathbb{T}^{m}$ is semialgebraic. Hence $\mathcal{J}$ is semialgebraic, and $\mathcal{D}$ is definable in $\mathbb{R}_{0}$. To prove (d), suppose additionally that $P\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$, and $b_{1},\ldots,b_{m}$ satisfy either (i) or (ii). Then $\mathcal{D}$ is semialgebraic, and by Theorems 5 and 6, a representation of $\mathcal{X}$ (and hence $\mathcal{D}$) can be effectively computed.
Due to the structure of $J$, for every $t\geq 0$ and $x\in\mathbb{R}^{d}$, $\det(e^{Jt})=1$ and $\bigl{\|}e^{Jt}x\bigr{\|}=\bigl{\|}e^{-Jt}x\bigr{\|}=\|x\|$, which implies that $\bigl{\|}e^{Jt}\bigr{\|}=\bigl{\|}e^{-Jt}\bigr{\|}=1$. Choose $b=\|P^{-1}\|\cdot\|P\|$. We have that for every $t\geq 0$, $\det(e^{Dt})=\det(P^{-1})\det(e^{Jt})\det(P)=1$, $\bigl{\|}e^{Dt}\bigr{\|}\leq\|P^{-1}\|\cdot\bigl{\|}e^{Jt}\bigr{\|}\cdot\|P\|\leq b$, and similarly $\bigl{\|}e^{-Dt}\bigr{\|}\leq b$. Statement (b) then follows from the continuity of the determinant, the matrix norm, and the matrix inverse.
Finally, by Theorem 3, for every non-empty open subset $O$ of $X$ there exist infinitely many $n\in\mathbb{N}$ such that $(e^{\bm{i}b_{1}n},\ldots,e^{\bm{i}b_{m}n})\in O$. Statement (c) then follows from the fact that $f$ is a homeomorphism. ∎
Corollary 3.
Let $s\in\mathbb{R}^{d}$ and $O$ be an open subset of $\mathcal{D}\cdot s$. There exist infinitely many $n\in\mathbb{N}$ such that $e^{Dt}s\in O$.
Proof.
Apply Lemma 2 (c) with the open subset $\{A\in\mathcal{D}\colon As\in O\}$ of $\mathcal{D}$. ∎
The corollary above states that every point in $\mathcal{D}\cdot s$ is an $\omega$-limit point of the trajectory $(e^{Dt}s)_{t\geq 0}$. We will also need the following lemma about the orbit of an open set in a dynamical system defined by a rotation matrix.
Lemma 4.
Let $O\subseteq\mathbb{R}^{d}$ be open and $H=\{X\cdot O\mid X\in\mathcal{D}\}$. For every $y\in H$ there exists $\varepsilon>0$ such that $e^{Dn}\cdot O\supset B(y,\varepsilon)$ for infinitely many $n\in\mathbb{N}$.
Proof.
Consider $y\in H$. By definition of $H$, there exist $A\in\mathcal{D}$ and $z\in O$ such that $y=Az$. Let $\varepsilon_{1},\varepsilon_{2}>0$ be such that $O\supseteq B(z,\varepsilon_{1}+\varepsilon_{2})$. By Lemma 2 (b), $A$ is invertible and hence $\|A\|>0$. Let $\delta_{1},\kappa>0$ be such that for all $C$ with $\|C-A\|<\delta_{1}$ we have $\|C\|>\kappa$. Further let $\delta_{2}=\kappa\varepsilon_{2}/\varepsilon_{1}$ and $\delta=\min\{\delta_{1},\delta_{2}\}$. Consider $F\in\mathcal{D}$ such that $\|F-A\|<\delta$. We have that
$\|Fz-y\|=\|(F-A)z\|\leq\|F-A\|\|z\|<\delta\varepsilon_{1}<\kappa\varepsilon_{2}.$
Hence
$F\cdot O\supseteq F\cdot B(z,\varepsilon_{1})+F\cdot B(\varepsilon_{2})\supseteq Fz+B(\kappa\varepsilon_{2})+F\cdot B(\varepsilon_{1})\supseteq y+F\cdot B(\varepsilon_{1}).$
It remains to choose $\varepsilon=\varepsilon_{1}/\|F\|$, so that $F\cdot B(\varepsilon_{1})\supseteq B(\varepsilon)$. ∎
5.2. Dynamics of scaling matrices
Now consider a scaling matrix $C\in\mathbb{R}^{d\times d}$ and $T\subseteq\mathbb{R}^{d}$ definable in an o-minimal expansion of $\mathbb{R}_{\exp}$. We will work with the sequence of sets $(e^{Ct}\cdot T)_{t\geq 0}$. Recall the definitions of Section 4.
Lemma 5.
The sequence $(e^{Ct}\cdot T)_{t\geq 0}$ converges to a limit shape $L$. In case $C\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ and $T$ is semialgebraic, $L$ is semialgebraic with an effectively computable representation.
Proof.
The sequence $(e^{Ct}\cdot T)_{t\geq 0}$ is definable in the same o-minimal expansion of $\mathbb{R}_{\exp}$ in which $T$ is definable. The first claim follows from Theorem 3. The second claim follows from Lemma 5 and the definition of the matrix exponential. ∎
5.3. Constructing decompositions
We now show that every $M$ has a decomposition. We also collect all properties of decompositions and effectiveness results that we will need.
Lemma 6.
Every $M\in\mathbb{R}^{d\times d}$ has a decomposition $(C,D)$ with the following properties.
  • (a)
    If $M=\operatorname{diag}(\Lambda(a_{1},b_{1}),\ldots,\Lambda(a_{m},b_{m}),a_{m+1},\ldots,a_{l})$, where $a_{i},b_{i}\in\mathbb{R}$ for all $i$, then
    $C$ $=\operatorname{diag}(\Lambda(a_{1},0),\ldots,\Lambda(a_{m},0),a_{m+1},\ldots,a_{l}),$
    $D$ $=\operatorname{diag}(\Lambda(0,b_{1}),\ldots,\Lambda(0,b_{m}),\mathbf{0}).$
  • (b)
    If $M\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$, then $C,D\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$. Moreover, (i) the closure $\mathcal{D}$ of $(e^{Dt})_{t\geq 0}$ is semialgebraic and effectively computable, and (ii) we can compute polynomials $p_{1,1},\ldots,p_{d,d}$ with real algebraic coefficients such that for all $t\geq 0$ and $1\leq i,j\leq d$,
    $(e^{-Ct})_{i,j}=p_{i,j}(t,e^{-a_{1}t},\ldots,e^{-a_{l}t})$
    where $a_{1},\ldots,a_{l}$ are the eigenvalues of $C$.
Proof.
Let $P,J$ be such that $J$ is in real Jordan form and $M=P^{-1}JP$. By permuting the blocks if necessary, we can assume that for some $m\geq 0$ and $l>0$, $J=\operatorname{diag}(J_{1},\ldots,J_{l})$, each $J_{i}$ is a real Jordan block, $J_{i}$ has a pair of non-real eigenvalues for $1\leq i\leq m$, and $J_{i}$ has a single real eigenvalue for $i>m$. Let $a_{i}$ for $1\leq i\leq l$ and $b_{i}$ for $1\leq i\leq m$ be such that
  • (i)
    for $i\leq m$, $J_{i}=\operatorname{diag}(\Lambda(a_{i},b_{i}),\ldots,\Lambda(a_{i},b_{i}))+N_{i}$ where $N_{i}$ is a nilpotent matrix, and
  • (ii)
    for $m<i\leq l$, $J_{i}=a_{i}I+N_{i}$ where $N_{i}$ is a nilpotent matrix.
We give a decomposition $(\widetilde{C}_{i},\widetilde{D}_{i})$ of $J_{i}$ for all $i$, which yields the decomposition $(\widetilde{C},\widetilde{D})$ of $J$ with $\widetilde{C}=\operatorname{diag}(\widetilde{C}_{1},\ldots,\widetilde{C}_{l})$ and $\widetilde{D}=\operatorname{diag}(\widetilde{D}_{1},\ldots,\widetilde{D}_{l})$; Thereafter we simply define $C=P^{-1}\widetilde{C}P$ and $D=P^{-1}\widetilde{D}P$. For $i\leq m$, let $\Gamma_{i}=\Lambda(a_{i},b_{i})-a_{i}I$, $D_{i}=\operatorname{diag}(\Gamma_{i},\ldots,\Gamma_{i})$, and $C_{i}=J_{i}-D_{i}$. For $m<i\leq l$, let $D_{i}=\mathbf{0}$ and $C_{i}=J_{i}$. It can be verified directly that $(C_{i},D_{i})$ is a decomposition of $J_{i}$. That this decomposition satisfies (a) is immediate from the construction.
To prove (b), let $(C,D)$ be the decomposition of $M$ constructed above. Note that $b_{1},\ldots,b_{m}\in\mathbb{R}\cap\overline{\mathbb{Q}}$, and $a_{1},\ldots,a_{l}$ are exactly the eigenvalues of $C$. Hence by Lemma 2, the closure of $(e^{Dt})_{t\geq 0}$ is semialgebraic and effectively computable. To prove (ii), first observe that
$e^{-Ct}=P^{-1}\operatorname{diag}(e^{-\widetilde{C}_{1}t},\ldots,e^{-\widetilde{C}_{l}t})P$
for all $t\geq 0$. Hence it suffices to consider the entries of each $e^{-\widetilde{C}_{i}t}$ separately. Recall that
(5) $e^{-\widetilde{C}_{i}t}=\begin{bmatrix}A_{i}(t)&tA_{i}(t)&\cdots&\frac{t^{\sigma}_{i}}{\sigma_{i}!}A_{i}(t)\\ &A_{i}(t)&\ddots&\vdots\\ &&\ddots&tA_{i}(t)\\ &&&A_{i}(t)\end{bmatrix}$
where $\sigma_{i}$ is the multiplicity of $-C_{i}$ and $A_{i}(t)=\operatorname{diag}(e^{-a_{i}t},e^{-a_{i}t})$ if $1\leq i\leq m$ and $A_{i}(t)=e^{-a_{i}t}$ otherwise. We see that each entry is polynomial in $t$ and $e^{-a_{i}t}$. ∎
Using the matrix logarithm, a discrete-time linear dynamical system can be embedded into a continuous-time linear dynamical system. This will allow us to transfer our solutions to the continuous-time variants of the Robust Safety Problem and the Pseudo-Reachability Problem to discrete-time systems. The next lemma summarises this embedding and the resulting decomposition.
Lemma 7.
Suppose $Q\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ is invertible. Then we can compute matrices $M,C,D$, a semialgebraic set $\mathcal{D}$, polynomials $p_{1,1},\ldots,p_{d,d}$ with real algebraic coefficients, and $\rho_{1},\ldots,\rho_{l}\in\mathbb{R}_{>0}\cap\overline{\mathbb{Q}}$ with the following properties.
  • (a)
    $(C,D)$ is a decomposition of $M$, and for all $n\in\mathbb{N}$, $Q^{n}=e^{Mn}$.
  • (b)
    $\mathcal{D}$ is the closure of $(e^{Dt})_{t\geq 0}$.
  • (c)
    For every $1\leq i,j\leq d$ and $t\geq 0$, $(e^{-Ct})_{i,j}=p_{i,j}(t,\rho_{1}^{t},\ldots,\rho_{l}^{t}).$
Proof.
We first compute $m,l\geq 0$ and $P,J\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ such that $J$ is in real Jordan form, $Q=P^{-1}JP$ and $J=\operatorname{diag}(J_{1},\ldots,J_{l})$, where for all $i$, $J_{i}=\operatorname{diag}(\Lambda_{i},\ldots,\Lambda_{i})+N_{i}$, each $N_{i}$ is nilpotent, $\Lambda_{i}=\rho_{i}R(\theta_{i})$ if $1\leq i\leq m$, and $\Lambda_{i}=[\rho_{i}]$ if $m<i\leq l$. We have that $\rho_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ and, by invertibility, $\rho_{i}>0$ for all $i$. Write $\log(\Lambda_{i})$ as $\Theta_{i}+\Gamma_{i}$, where
  • (i)
    for $1\leq i\leq m$, $\Theta_{i}\coloneqq\Lambda(0,b_{i})$ for some $b_{i}\in\{\operatorname{Log}(z)/\bm{i}\mid z\in\mathbb{T}\cap\overline{\mathbb{Q}}\}$ and $\Gamma_{i}=\Lambda(\log(\rho_{i}),0)$, and
  • (ii)
    for $i>m$, $\Theta_{i}=[0]$ and $\Gamma_{i}=[\log(\rho_{i})]$.
Next, for $1\leq i\leq l$, compute $\log(J_{i})$ and let $C_{i},D_{i}$ be such that $\log(J_{i})=C_{i}+D_{i}$, $C_{i}$ is nilpotent and has real algebraic entries, and $D_{i}=\operatorname{diag}(\log(\Lambda_{i}),\ldots,\log(\Lambda_{i}))$. For $1\leq i\leq m$, define $\widetilde{C}_{i}=\operatorname{diag}(\Gamma_{i},\ldots,\Gamma_{i})+C_{i}$ and $\widetilde{D}_{i}=\operatorname{diag}(\Theta_{i},\ldots,\Theta_{i})$, so that $\widetilde{C}_{i},\widetilde{D}_{i},J_{i}$ have the same dimensions. Finally, let $\widetilde{C}=\operatorname{diag}(\widetilde{C}_{1},\ldots,\widetilde{C}_{l})$, $\widetilde{D}=\operatorname{diag}(\widetilde{D}_{1},\ldots,\widetilde{D}_{l})$, $C=P^{-1}\widetilde{C}P$, $D=P^{-1}\widetilde{D}P$, and $M=C+D$. It can be verified directly that each $(\widetilde{C}_{i},\widetilde{D}_{i})$ is a decomposition of $\log(J_{i})$ and hence $(\widetilde{C},\widetilde{D})$ is a decomposition of $\log(J)$ and $(C,D)$ is a decomposition of $M$. Consequently,
$Q^{n}=P^{-1}e^{\log(J)n}P=P^{-1}e^{\widetilde{C}n}e^{\widetilde{D}n}P=P^{-1}e^{\widetilde{C}n}P\cdot P^{-1}e^{\widetilde{D}n}P=e^{Cn}e^{Dn}.$
We have so far shown (a). Since $b_{i}\in\{\operatorname{Log}(z)/\bm{i}\mid z\in\mathbb{T}\cap\overline{\mathbb{Q}}\}$ for all $i$, by Lemma 2, the closure of $(e^{Dt})_{t\geq 0}$ is semialgebraic and effectively computable. To prove (c), it suffices to consider the entries of $e^{-\widetilde{C}_{i}t}$ for $1\leq i\leq l$. We have that (5) holds with $A_{i}(t)=\operatorname{diag}(\rho_{i}^{-t},\rho_{i}^{-t})$ for $1\leq i\leq m$ and $A_{i}(t)=[\rho_{i}^{-t}]$ for $i>m$. That is, the entries of $e^{-\widetilde{C}_{i}t}$ are polynomial in $t$ and $\rho_{i}^{-t}$. ∎
6. The Robust Safety Problem
In this section we study the Robust Safety Problem for both discrete-time and continuous-time systems. Recall that in the continuous-time case we have to decide, given $M$ and semialgebraic $S,T$, whether there exists $\varepsilon>0$ such that the sequence $(e^{Mt}\cdot B(S,\varepsilon))_{t\geq 0}$ avoids $T$. We will be applying the decomposition method. Consider a decomposition $(C,D)$ of $M$, and recall that $e^{Mt}=e^{Ct}e^{Dt}$ for all $t\geq 0$. We have that $e^{Mt}\cdot B(S,\varepsilon)\cap T=\emptyset$ if and only if $e^{Dt}\cdot B(S,\varepsilon)\cap e^{-Ct}\cdot T=\emptyset$. We will separately analyse the sequences of sets $(e^{Dt}\cdot(B(S,\varepsilon))_{t\geq 0}$ and $(e^{-Ct}\cdot T)_{t\geq 0}$, under the assumption that $S$ is bounded, using the theory we developed in Sections 5.1 and 5.2. First, we give a topological result, leaving effectiveness issues for later. The constructions of the next theorem are illustrated in Figure 1. Observe that if $T$ is definable in an o-minimal expansion $\mathbb{S}$ of $\mathbb{R}_{\exp}$, then the sequence $(e^{-Ct})_{t\geq 0}$ is definable in $\mathbb{S}$. From Lemma 5 it then follows that $(e^{-Ct})_{t\geq 0}$ converges to a limit shape.
Theorem 1.
Let $M\in\mathbb{R}^{d\times d}$, $S\subseteq\mathbb{R}^{d}$ be bounded, and $T\subseteq\mathbb{R}^{d}$. Suppose $S,T$ are definable in an o-minimal expansion of $\mathbb{R}_{\exp}$. Let $(C,D)$ be a decomposition of $M$, $L$ be the limit shape of $(e^{-Ct}\cdot T)_{t\geq 0}$, and $\mathcal{D}$ be the closure of $(e^{Dt})_{t\geq 0}$. Define $H(\varepsilon)=\mathcal{D}\cdot B(S,\varepsilon)$ for $\varepsilon\geq 0$, and
(6) $\mu=\inf\{\varepsilon\geq 0\colon L\textrm{ intersects }\>H(\varepsilon)\}.$
We have the following.
  • (a)
    For every $\varepsilon>\mu$, there exist infinitely many $n\in\mathbb{N}$ such that
    $e^{Mn}\cdot B(S,\varepsilon)\>\textrm{ intersects }\>T.$
  • (b)
    For every $0\leq\varepsilon<\mu$, there exists $N\in\mathbb{N}$ such that $(e^{-Ct}\cdot T)_{t\geq N}$ avoids $H(\varepsilon)$ and
    $(e^{Mt}\cdot B(S,\varepsilon))_{t\geq N}\>\textrm{ avoids }\>T.$
Proof.
Applying the decomposition method,
(7) $e^{Mt}\cdot B(S,\varepsilon)\textrm{ intersects }T\Leftrightarrow e^{Dt}\cdot B(S,\varepsilon)\textrm{ intersects }e^{-Ct}\cdot T$
for all $t\geq 0$. Observe that for every $\varepsilon>0$, $B(S,\varepsilon)$ is bounded and open. Intuitively, $e^{Dt}\cdot B(S,\varepsilon)$ is a bounded open ball that moves around (as $t$ continuously increases) in the bounded set $H(\varepsilon)$ in a uniformly recurrent manner: Every region of $H(\varepsilon)$ is visited infinitely often by Lemma 4. The sequence $(e^{-Ct}\cdot T)_{t\geq 0}$, on the other hand, converges to a limit shape $L$. This allows us to treat $e^{-Ct}\cdot T$ as stable for sufficiently large $t$. Hence in all cases except where $\varepsilon=\mu$, i.e. $H(\varepsilon)$ is not topologically separated from the limit shape $L$, we are able to determine whether there exist arbitrarily large values of  $t$ such that $e^{Dt}\cdot B(S,\varepsilon)$ intersects $e^{-Ct}\cdot T$. We next formalise this argument.
Since $e^{Dt}\in\mathcal{D}$, we have that $e^{Dt}\cdot B(S,\varepsilon)\subseteq H(\varepsilon)$ for all $\varepsilon,t\geq 0$. To prove (a), consider $\varepsilon>\mu$. By construction of $\mu$, $L$ intersects $H(\varepsilon)$. Choose $y\in L\cap H(\varepsilon)$. Applying Lemma 4, let $\varepsilon_{1}>0$ be such that there exist infinitely many $n\in\mathbb{N}$ satisfying $e^{Dn}\cdot B(S,\varepsilon)\supset B(y,\varepsilon_{1})$. Since $y\in L$, by definition of the limit shape there exists $N$ such that for all $t\geq N$, $B(y,\varepsilon_{1})$ intersects $e^{-Ct}\cdot T$. It follows that there exist infinitely many $n\in\mathbb{N}$ such that $e^{Dn}\cdot B(S,\varepsilon)$ intersects $e^{-Cn}\cdot T$. Finally, apply (7).
We next prove (b). If $\mu=0$, then the statement if vacuously true. Suppose $\mu>0$, and consider $0\leq\varepsilon<\mu$. Invoking Lemma 2 (b), let $b$ be such that $\|A^{-1}\|<b$ for all $A\in\mathcal{D}$. Let $X=\operatorname{Cl}(H(\varepsilon))$, $\varepsilon_{1}\in(\varepsilon,\mu)$, and $\gamma=(\varepsilon_{1}-\varepsilon)/b>0$. Observe that $X$ is compact. We will show that $B(L,\gamma)$ is disjoint from $X$. To do this, we first argue that
(8) $B(H(\varepsilon),\gamma)\subseteq H(\varepsilon_{1}).$
For every $A\in\mathcal{D}$ and $x\in B(\gamma)$ we have that $\|A^{-1}x\|\leq\|A^{-1}\|\cdot\|x\|<\varepsilon_{1}-\varepsilon$. Consider $A(y+z)+x\in B(H(\varepsilon),\gamma)$ with $A\in\mathcal{D}$, $y\in S$, $z\in B(\varepsilon)$, and $x\in B(\gamma)$. We have
$A(y+z)+x=A(y+(z+A^{-1}x))\in A\cdot B(S,\varepsilon_{1})\in H(\varepsilon_{1})$
which proves (8). By construction of $\mu$, $\operatorname{Cl}(H(\varepsilon_{1}))$ is disjoint from $L$. Hence $B(X,\gamma)$, which satisfies
$B(X,\gamma)\subseteq\operatorname{Cl}(B(H(\varepsilon),\gamma))\subseteq\operatorname{Cl}(H(\varepsilon_{1})),$
is disjoint from $L$. It follows that $B(L,\gamma)$ is disjoint from $X$.
By definition of the limit shape there exists $N\in\mathbb{N}$ such that
$(e^{-Ct}\cdot T)\cap X\subseteq B(L\cap X,\gamma)$
for all $t\geq N$. Because $B(L,\gamma)$ is disjoint from $X$, for all $t\geq N$, $(e^{-Ct}\cdot T)\cap X=\emptyset$. Since $e^{Dt}\cdot B(S,\varepsilon)\subseteq H(\varepsilon)\subseteq X$ for all $t\geq 0$, we have that $e^{-Ct}\cdot T$ does not intersect $e^{Dt}\cdot B(S,\varepsilon)$ for all $t\geq N$. ∎
$L$$e^{-Ct}\cdot T$$S$$H(0)$$H(\mu)$
Figure 1. The constructions of Theorem 1. Here $\mathcal{D}$ is the group of all $2\times 2$ rotation matrices, the larger annulus is $H(\mu)$ and the smaller one is $H(0)=\mathcal{D}\cdot S\subseteq H(\mu)$. The sequence ($e^{-Ct}\cdot T)_{t\geq 0}$ converges to $L$.
Note that the theorem above makes no claims about the set $\{t\geq 0\colon e^{Mt}\cdot B(S,\mu)\textrm{ intersects }T\}$. This is unsurprising for the following reasons.
  • (i)
    The Safety Problem is equivalent to the problem of deciding whether $e^{Mt}\cdot B(S,0)$ intersects $T$ for some $t\geq 0$.
  • (ii)
    The hard cases of the Safety Problem considered in, whose decidability would lead to major breakthroughs in Diophantine approximation, arise when $M\in\mathbb{Q}^{d\times d}$, $S,T$ are semialgebraic, $S$ is a singleton, and $\mu=0$.
Therefore, understanding the structure of $\{t\geq 0\colon e^{Mt}\cdot B(S,\varepsilon)\textrm{ intersects }T\}$ for $\varepsilon=\mu=0$ would likely at least require sophisticated arguments in Diophantine approximation. Nevertheless, we are able to show decidability of the Robust Safety Problem. We will need a small lemma.
Lemma 2.
Let $M\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$, $\varphi\in\mathcal{L}_{or}$ and $N\in\mathbb{N}$. The set of all $t\in[0,N]$ such that $\varphi((e^{Mt})_{1,1},\ldots,(e^{Mt})_{d,d},t)$ holds is definable in $\mathbb{R}_{\exp,\,\operatorname{bt}}$ without parameters.
Proof.
Let $a_{1}+\bm{i}b_{1},\ldots,a_{d}+\bm{i}b_{d}$ be the eigenvalues of $M$ with $a_{i},b_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$. Further let $B\in\mathbb{Q}$ be such that $|b_{i}|<B$ for all $i$, and define $c_{i}=b_{i}/(2BN))$. As shown in Section 2.5, each entry of $e^{Mt}$ is polynomial in $e^{a_{i}t}$, $\sin(b_{i}t)$, $\cos(b_{i}t)$ for $1\leq i\leq d$. Recall the definitions of the functions $\widetilde{\cos},\widetilde{\sin}$, and that both $\sin(b_{i}t)$ and $\cos(b_{i}t)$ are polynomial in $\sin(c_{i}t)$ and $\cos(c_{i}t)$. We can therefore construct, by performing the linear change of variables $u=t/(2BN)$, a formula $\psi\in\mathcal{L}_{or}$ such that
$\exists t\in[0,N]\colon\varphi((e^{Mt})_{1,1},\ldots,(e^{Mt})_{d,d},t)\Leftrightarrow\exists u\in[0,1/2]\colon\psi(e^{a^{\prime}_{1}u},\sin(b^{\prime}_{1}u),\cos(b^{\prime}_{1}u),\ldots,e^{a^{\prime}_{d}u},\sin(b^{\prime}_{d}u),\cos(b^{\prime}_{d}u))$
where $a^{\prime}_{i}=a_{i}\cdot 2BN$ and $b^{\prime}_{i}=b_{i}/(2BN)\in(0,1)$ for all $i$. Since $b^{\prime}_{i}u\in(-1,1)$ for all $u\in[0,1/2]$, we can replace $\cos,\sin$ with $\widetilde{\cos},\widetilde{\sin}$, respectively, in the right-hand side of the equivalence above. ∎
Proof of Theorem 1.
Let $\langle M,S,T\rangle$ be an instance of the Robust Safety Problem in dimension $d$ with bounded $S$. Applying Lemma 6, compute a decomposition $(C,D)$ of $M$, the semialgebraic set $\mathcal{D}$, real algebraic $a_{1},\ldots,a_{l}>0$, and the formula $\varphi\in\mathcal{L}_{or}$ such that for all $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$,
$x\in e^{-Ct}\cdot T\Leftrightarrow\varphi(x_{1},\ldots,x_{d},t,e^{-a_{1}t},\ldots,e^{-a_{l}t}).$
Recall that $\mathcal{D}$ is the closure of $(e^{Dt})_{t\geq 0}$. Applying Lemma 5, the limit shape $L$ of $(e^{-Ct}\cdot T)_{t\geq 0}$ is semialgebraic and effectively computable. Next, define $\mu$ and $H(\varepsilon)$ as in Theorem 1. If $L=\emptyset$, then $\mu=+\infty$. Otherwise, $\mu\in\mathbb{R}_{\geq 0}\cap\overline{\mathbb{Q}}$, and can be effectively computed from $L$ and $\mathcal{D}$. If $\mu=0$, then $\langle M,S,T\rangle$ is a negative instance: By Theorem 1 (a), for every $\varepsilon>0$ there exist infinitely many $n\in\mathbb{N}$ such that $e^{Mn}\cdot B(S,\varepsilon)$ intersects $T$.
Suppose therefore that $\mu>0$. Choose $\varepsilon_{1}\in(0,\mu)\cap\mathbb{Q}$. We will compute $N\in\mathbb{N}$ such that $(e^{Mt}\cdot B(S,\varepsilon_{1}))_{t\geq N}$ avoids $T$. Construct a formula $\psi\in\mathcal{L}_{or}$ such that for all $\varepsilon>0$ and $t\geq 0$,
$H(\varepsilon)\textrm{ intersects }e^{-Ct}\cdot T\Leftrightarrow\psi(\varepsilon,t,e^{a_{1}t},\ldots,e^{a_{l}t}).$
By Theorem 1 (b), the statement $\psi(\varepsilon_{1},t,e^{a_{1}t},\ldots,e^{a_{l}t})$ evaluates to false for all sufficiently large $t$. It remains to apply Lemma 4 (c) to compute $N\in\mathbb{N}$ such that $\psi(\varepsilon_{1},t,e^{-a_{1}t},\ldots,e^{a_{l}t})$ does not hold for all $t\geq N$, and observe that $\lnot\psi(\varepsilon_{1},t,e^{a_{1}t},\ldots,e^{a_{l}t})$ implies that $e^{Mt}\cdot B(S,\varepsilon_{1})$ does not intersect $T$.
Having shown that $e^{Mt}\cdot B(S,\varepsilon_{1})$ does not intersect $T$ for all $t\geq N$, it remains to understand what happens for $t\in[0,N)$. By Lemma 2, assuming Schanuel’s conjecture we can decide whether
$\exists t\in[0,N]\colon e^{Mt}\cdot\operatorname{Cl}(S)\textrm{ intersects }\operatorname{Cl}(T).$
If the sentence above is true, then for every $\varepsilon>0$ there exists $t\in[0,N)$ such that $e^{Mt}\cdot B(S,\varepsilon)$ intersects $T$. Hence $\langle M,S,T\rangle$ is a negative instance. Otherwise, there exists $\varepsilon_{2}\in\mathbb{Q}_{>0}$ such that for all $t\in[0,N)$, $e^{Mt}\cdot B(S,\varepsilon_{2})$ does not intersect $T$. Such $\varepsilon_{2}$ can be computed as follows. Denote the eigenvalues of $M$ by $u_{1}+\bm{i}v_{1},\ldots,u_{d}+\bm{i}v_{d}$, where $u_{i},v_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$, and recall that the entries of $e^{Mt}$ are polynomial in $t$ and $e^{u_{i}t},\cos(v_{i}t),\sin(v_{i}t)$ for $1\leq i\leq d$. Using Taylor expansions of the exponential and trigonometric functions, for every $\kappa>0$ we can compute $\varphi_{\kappa}\in\mathcal{L}_{or}$ such that for all $0\leq t<N$,
$e^{Mt}\in\underbrace{\{X\in\mathbb{R}^{d\times d}\colon\varphi_{\kappa}(t,X_{1,1},\ldots,X_{d,d})\}}_{\coloneqq X_{\kappa}(t)}\subseteq B(e^{Mt},\kappa).$
Because $\{e^{Mt}\cdot\operatorname{Cl}(S)\colon 0\leq t\leq N\}$ and $\operatorname{Cl}(T)$ are topologically separated by assumption, there exist $\kappa^{*},\varepsilon^{*}>\in\mathbb{Q}_{>0}$ such that $X_{\kappa}(t)\cdot B(S,\varepsilon)$ does not intersect $T$ for all $0\leq t\leq N$, $0\leq\kappa\leq\kappa^{*}$ and $0\leq\varepsilon\leq\varepsilon^{*}$. Such $\kappa^{*}$ and $\varepsilon^{*}$ can be determined using a trial-and-error method and decision procedures for semialgebraic sets. We can then choose $\varepsilon_{2}=\varepsilon^{*}$. Finally, let $\tilde{\varepsilon}=\min\{\varepsilon_{1},\varepsilon_{2}\}\in\mathbb{Q}_{>0}$. We have that the sequence $(e^{Mt}\cdot B(S,\tilde{\varepsilon}))_{t\geq 0}$ avoids $T$. ∎
6.1. The Discrete-Time Robust Safety Problem
We now prove Theorem 2. Let $\langle Q,S,T\rangle$ be an instance of the Discrete-Time Robust Safety Problem in dimension $d$ with bounded $S$. First assume $Q$ is invertible; We will handle the general case later. Compute the constructions of Lemma 7 for $Q$, recalling that $Q^{n}=e^{Mn}$ for all $n\in\mathbb{N}$. Applying the decomposition method,
$Q^{n}\cdot B(S,\varepsilon)\textrm{ intersects }T\Leftrightarrow e^{Dn}\cdot B(S,\varepsilon)\textrm{ intersects }e^{-Cn}\cdot T$
for all $\varepsilon>0$ and $n\in\mathbb{N}$. Applying Lemma 7 (f), let $\varphi\in\mathcal{L}_{or}$ be such that for all $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$,
$x\in e^{-Ct}\cdot T\Leftrightarrow\varphi(x_{1},\ldots,x_{l},\rho_{1}^{t},\ldots,\rho_{l}^{t}).$
By Theorem 3, $(e^{-Ct})_{t\geq 0}$ converges to a limit shape $L$. By Lemma 5, $L$ is semialgebraic and effectively computable. We can now apply the same analysis as in the proof of Theorem 1. Define $H(\varepsilon)=\mathcal{D}\cdot B(S,\varepsilon)$, and compute
$\mu=\inf\{\varepsilon>0\colon L\textrm{ intersects }H(\varepsilon)\},\quad\mu\in\mathbb{R}_{>0}\cup\{+\infty\}.$
If $\mu=0$, then $\langle Q,S,T\rangle$ is a negative instance: By Lemma 4 (a), for every $\varepsilon>0$ there exist infinitely many $n\in\mathbb{N}$ such that $Q^{n}\cdot B(S,\varepsilon)$ intersects $T$.
Suppose $\mu>0$. Let $\varepsilon_{1}\in(0,\mu)\cap\mathbb{Q}$. We will compute $N\in\mathbb{N}$ such that $(Q^{n}\cdot B(S,\varepsilon_{1}))_{n\geq d}$ avoids $T$. Using the semialgebraic set $\mathcal{D}$ and the formula $\varphi$, construct $\psi\in\mathcal{L}_{or}$ such that for all $\varepsilon,t\geq 0$,
$H(\varepsilon)\textrm{ intersects }e^{-Ct}\cdot T\Leftrightarrow\psi(\varepsilon,t,\rho_{1}^{t},\ldots,\rho_{l}^{t}).$
By Theorem 1 (b), the statement $\psi(\varepsilon_{1},t,\rho_{1}^{t},\ldots,\rho_{l}^{t})$ evaluates to false for all sufficiently large $t$. Applying Lemma 4 (c) we can compute $N\in\mathbb{N}$ such that $\psi(\varepsilon_{1},t,\rho_{1}^{t},\ldots,\rho_{l}^{t})$ does not hold for all $t\geq N$. Finally, observe that $\lnot\psi(\varepsilon_{1},n,\rho_{1}^{n},\ldots,\rho_{l}^{n})$ implies that $Q^{n}\cdot B(S,\varepsilon_{1})$ does not intersect $T$.
Having computed $N$, let
$\varepsilon_{2}=\sup\{\varepsilon>0\colon Q^{n}\cdot B(S,\varepsilon)\cap T=\emptyset\textrm{ for }n=0,\ldots,N-1\}.$
If $\varepsilon_{2}=0$, then again $\langle Q,S,T\rangle$ is a negative instance: for every $\varepsilon>0$, there exists $0\leq n<N$ such that $Q^{n}\cdot B(S,\varepsilon)$ intersects $T$. Otherwise, define $\tilde{\varepsilon}=\min\{\varepsilon_{1},\varepsilon_{2}\}/2$. We have that $(Q^{n}\cdot B(S,\tilde{\varepsilon}))_{n\in\mathbb{N}}$ avoids $T$.
Proof of Theorem 2.
Let $\langle M,S,T\rangle$ be an instance of the Discrete-Time Robust Safety Problem. Write $M=P^{-1}JP$, where $J=\operatorname{diag}(J_{1},J_{2})$ is in real Jordan form, $J_{1}\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d_{1}\times d_{1}}$ is invertible, and $J_{2}$ is nilpotent. Observe that for all $\varepsilon>0$, $(M^{n}\cdot B(S,\varepsilon))_{n\in\mathbb{N}}$ avoids $T$ if and only if $(J^{n}P\cdot B(S,\varepsilon))_{n\in\mathbb{N}}$ avoids $P\cdot T$. Let $\pi\colon\mathbb{R}^{d}\to\mathbb{R}^{d_{1}}$ be the projection operation onto the first $d_{1}$ coordinates, $S_{1}=\pi(P\cdot S)$, $O=P\cdot B(S,\varepsilon)$, $O_{1}=\pi(O)$, and $T_{1}$ be the set of all $x\in T$ whose last $d-d_{1}$ coordinates are zero. Since the last $d-d_{1}$ coordinates of $J^{n}x$ are identically zero for all $x\in\mathbb{R}^{d}$ and $n\geq d$, we have that for all $\varepsilon>0$, $(J^{n}P\cdot B(S,\varepsilon))_{n\in\mathbb{N}}$ avoids $P\cdot T$ if and only if
  • (a)
    $J^{n}P\cdot B(S,\varepsilon)$ does not intersect $P\cdot T$ for $n=0,\ldots,d-1$, and
  • (b)
    $(J_{1}^{n}(S_{1}+\varepsilon O_{1}))_{n\in\mathbb{N},\,n\geq d}$ avoids $T_{1}$, which is the case if and only if
    (9) $(J_{1}^{n}(J_{1}^{d}\cdot S_{1}+\varepsilon J_{1}^{d}\cdot O_{1}))_{n\in\mathbb{N}}\quad\textrm{ avoids }\quad T_{1}.$
Because $O_{1}$ is an open set containing the origin, we can compute $\kappa_{1},\kappa_{2}>0$ be such that
$B(\kappa_{1})\subseteq J_{1}^{d}\cdot O_{1}\subseteq B(\kappa_{2}).$
Since $J_{1}$ is invertible, using our analysis above on the instance $\langle J_{1},J_{1}^{d}\cdot S_{1},T_{1}\rangle$ we can decide whether there exists $\varepsilon>0$ such that
$(J_{1}^{n}(J_{1}^{d}\cdot S_{1}+B(\varepsilon))_{n\in\mathbb{N}}\quad\textrm{ avoids }\quad T_{1},$
and compute $\varepsilon\in\mathbb{Q}_{>0}$ with the property above in case it exists. If no such $\varepsilon$ exists, then $\langle M,S,T\rangle$ is a negative instance: For every $\varepsilon>0$, there exists $n$ such that $J_{1}^{n}(J_{1}^{d}\cdot S_{1}+B(\varepsilon)$ and hence $J_{1}^{n}\cdot(J_{1}^{d}\cdot S_{1}+\frac{\varepsilon}{\kappa_{2}}J_{1}^{d}\cdot O_{1})$ intersect $T$.
Now suppose we computed $\varepsilon_{1}\in\mathbb{Q}_{>0}$ such that $\varepsilon=\varepsilon_{1}$ satisfies (9). Define
$\varepsilon_{2}=\sup\{x=\varepsilon\geq 0\colon J^{n}(P\cdot S+\varepsilon P\cdot O)\textrm{ avoids $P\cdot T$ for }n=0,\ldots,N-1\},$
and $\tilde{\varepsilon}=\min\{\varepsilon_{1},\varepsilon_{2}\}/2$. If $\varepsilon_{2}\leq 0$, then $\langle M,S,T\rangle$ is a negative instance. Otherwise, $\tilde{\varepsilon}>0$ and $(M^{n}\cdot B(S,\tilde{\varepsilon}))_{n\in\mathbb{N}}$ avoids $T$. ∎
6.2. Connection to o-minimal invariants
Invariants are well-known objects in dynamical systems theory and ergodic theory. In formal verification, they are an important tool for certifying non-reachability. A set $I\subseteq\mathbb{R}^{d\times d}$ is an (inductive) invariant of a discrete-time LDS given by $M\in\mathbb{R}^{d\times d}$ if $M\cdot I\subseteq I$. Similarly, for a continuous-time LDS given by $M$, an invariant is a set $I$ such that $e^{Mt}\cdot I\subseteq I$ for all $t\geq 0$. Given a starting set $S$ and a set $T$ of unsafe configurations, if we can find an inductive invariant $I$ of $M$ such that $I\supseteq S$ and $I\cap T=\emptyset$, then we have a certificate that every trajectory starting in $S$ is safe, i.e. does not reach $T$. We say that such $I$ separates $S$ from $T$.
The Invariant Synthesis Problem for a class $\mathcal{I}$ is to decide, given $\langle M,S,T\rangle$, whether there exists an invariant $I\in\mathcal{I}$ of the LDS defined by $M$ that separates $S$ from $T$. This problem has been considered for various classes of invariants, including algebraic sets, semialgebraic sets, unions of subspaces, convex sets, and so on. We now discuss the connections between our work and (almagor2020cont-invariants, ; almagor2022discrete-invariants, ) which study o-minimal invariants, i.e. invariants definable in some expansion of $\mathbb{R}_{\exp}$. They only consider singleton $S$. Intuitively, if $\langle M,S,P\rangle$ is a positive instance of the Robust Safety Problem, then the sequence $(e^{Mt}\cdot S)_{t\geq 0}$ is topologically separated from $T$. This separation can be used to construct a non-trivial invariant.
Theorem 3.
Suppose $S$ is bounded and $\langle M,S,T\rangle$ is a positive instance of the Robust Safety Problem. Then there exists an invariant $I$ of the continuous-time LDS defined by $M$ separating $S$ from $T$ that is definable in $\mathbb{R}_{\exp,\,\operatorname{bt}}$.
Proof.
As in the proof of Theorem 1, compute a decomposition $(C,D)$ of $M$ and semialgebraic $\mathcal{D},L$ such that $\mathcal{D}$ is the closure of $(e^{Dt})_{t\geq 0}$ and $L$ be the limit shape of $(e^{-Ct}\cdot T)_{t\geq 0}$. Define $\mu$ as in (9). Because $\langle M,S,T\rangle$ is a negative instance, $\mu>0$. Let $0\leq\varepsilon<\mu$, and $H(\varepsilon)=\mathcal{D}\cdot B(S,\varepsilon)$. Following the proof of Theorem 1, we can compute $N\in\mathbb{N}$ such that $(e^{-Ct}\cdot T)_{t\geq N}$ avoids $H(\varepsilon)$. Let
$I=\{e^{Mt}\cdot S\mid 0\leq t<N\}\>\cup\>\{e^{Ct}\cdot\mathcal{D}\cdot S\mid t\geq N\}$
which is definable in $\mathbb{R}_{\exp,\,\operatorname{bt}}$ and disjoint from $T$. We will argue that $I$ is an invariant. Let $\tau\geq 0$. Because $e^{Dt}\in\mathcal{D}$ for all $t\geq 0$, the orbit $\{e^{Mt}\cdot S\mid t\geq 0\}$ is contained in $I$. Hence $e^{M\tau}\cdot\{e^{Mt}\cdot S\mid 0\leq t<N\}\subseteq I$. On the other hand,
$e^{M\tau}\cdot\{e^{Ct}\cdot\mathcal{D}\cdot S\mid t\geq N\}=\{e^{D\tau}e^{C\tau}e^{Ct}\cdot\mathcal{D}\cdot S\mid t\geq N\}.$
Because $C,D$ commute, $Ct_{1},Dt_{2}$ commute for every $t_{1},t_{2}\geq 0$. Moreover, $e^{Dt}\cdot\mathcal{D}=\mathcal{D}$ for all $t\geq 0$. Hence
$e^{M\tau}\cdot\{e^{Ct}\cdot\mathcal{D}\cdot S\mid t\geq N\}=\{e^{C(\tau+t)}\cdot\mathcal{D}\cdot S\mid t\geq N\}\subseteq I.$
We conclude that $e^{Mt}\cdot I\subseteq I$. ∎
The converse is not true: consider $M=\begin{bmatrix}0&-1&0\\ 1&0&0\\ 0&0&1\end{bmatrix}$ which satisfies $e^{Mt}=\begin{bmatrix}\cos(t)&-\sin(t)&0\\ \sin(t)&\cos(t)&0\\ 0&0&e^{t}\end{bmatrix}$. Let $S(r,h)=\{(x,y,h)\colon x^{2}+y^{2}=r^{2}\}$, $S=S(1,1)$, and $T=\bigcup_{h\geq 1}S(1+1/h,h)$. Then $e^{Mt}\cdot S=S(1,e^{t})$, $I=\bigcup_{h\geq 1}S(1,h)$ is an invariant separating $S$ from $T$, but $\langle M,S,T\rangle$ is a negative instance of the Robust Safety Problem.
An analogous result to Theorem 3 holds for discrete-time LDS. In this case, we can strengthen $\mathbb{R}_{\exp,\,\operatorname{bt}}$ to $\mathbb{R}_{\exp}$.
Theorem 4.
Suppose $S$ is bounded and $\langle M,S,T\rangle$ is a positive instance of the Discrete-Time Robust Safety Problem. Then there exists an invariant $I$ of the discrete-time LDS defined by $M$ that is definable in $\mathbb{R}_{\exp}$ and separates $S$ from $T$.
Proof.
Let $P,J\in\mathbb{R}^{d\times d}$ be such that $J$ is in real Jordan form, $M=P^{-1}JP$, and $J=\operatorname{diag}(J_{1},J_{2})$, where $J_{1}\in\mathbb{R}^{d_{1}\times d_{1}}$ is invertible and $J_{2}$ is nilpotent. Observe that $P^{-1}\cdot I$ is an invariant of the (discrete-time) LDS defined by $M$ separating $S$ from $T$ if and only if $I$ is an invariant of the LDS defined by $J$ separating $S$ from $T$. To construct such $I$, we will give $I_{1}$ that is an invariant of $J_{1}$ that separates $S_{1}\coloneqq\Pi(J^{d}\cdot S)$ from $T_{1}\coloneqq\Pi(T)$, where $\Pi\colon\mathbb{R}^{d}\to\mathbb{R}^{d_{1}}$ projects onto the first $d_{1}$ coordinates. Recalling that the last $d-d_{1}$ coordinates of $J^{n}x$ are identically zero for all $x$ and $n\geq d$, we then take
$I=\bigcup_{n=0}^{d-1}M^{n}\cdot S\cup\{(x_{1},\ldots,x_{d_{1}},0,\ldots,0)\mid(x_{1},\ldots,x_{d_{1}})\in I_{1}\}.$
Recall form the proof of Theorem 2 that $\langle J_{1},S_{1},T_{1}\rangle$ must be a positive instance of the Discrete-Time Robust Safety Problem. Compute $C,D,\mathcal{D},\mu$ as in the proof of Theorem 2. We must have $\mu>0$. Let $\varepsilon\in\mathbb{Q}_{>0}$ and $H(\varepsilon)=\mathcal{D}\cdot B(S_{1},\varepsilon)$. Recall that we can compute $N\in\mathbb{N}$ such that $(e^{-Ct}\cdot T_{1})_{t\geq N}$ avoids $H(\varepsilon)$. Let
$I=\bigcup_{n=0}^{N-1}M^{n}\cdot S\>\cup\>\{e^{Ct}\cdot\mathcal{D}\cdot S\mid t\geq N\}$
which is definable in $\mathbb{R}_{\exp}$ and separates $S$ from $T$. That $I$ is an invariant is shown in the same way as in the proof of Theorem 3: For every $\nu\in\mathbb{N}$, we have that
$M^{\nu}\cdot I=\bigcup_{n=\nu}^{\nu+N-1}M^{n}\cdot S\>\cup\>\{e^{C(t+\nu)}\cdot\mathcal{D}\cdot S\mid t\geq N\}.$
7. The Pseudo-Orbit Safety Problem
We now move on to the Pseudo-Orbit Safety Problem and its discrete-time variant. Our approach to showing decidability will be similar to the one used in Section 6, but we will need a few new ideas. Recall the definitions of the set-valued functions $\mathcal{R}$ and $\mathcal{P}$ from the Introduction. We have to decide, given $\langle M,S,T\rangle$, whether there exists $\varepsilon>0$ such that $(\mathcal{R}(M,S,B(\varepsilon),t))_{t\geq 0}$ avoids $T$. Let $(C,D)$ be a decomposition of $M$. Since $\mathcal{R}(M,S,B(\varepsilon),t)=e^{Mt}S+\varepsilon\cdot\mathcal{P}(M,B(1),t)$, using the decomposition method, this is equivalent to deciding whether there exists $\varepsilon>0$ such that for all $t\geq 0$,
$e^{Dt}\cdot S+\varepsilon e^{-Ct}\cdot\mathcal{P}(M,B(1),t)\>\textrm{ does notintersect }\>e^{-Ct}\cdot T.$
There are two major differences to the Robust Safety Problem. First, the sequence $(e^{-Ct}\cdot\mathcal{P}(M,B(1),t))_{t\geq 0}$ is not bounded, even if $S$ is. [^6] For comparison, in Section 6, we analysed sequences of the form $(e^{Dt}\cdot B(S,\varepsilon))_{t\geq 0}$, which can be written as $e^{Dt}\cdot B(S,\varepsilon)=e^{Dt}\cdot S+e^{Dt}\cdot B(\varepsilon)$. Here if $S$ is bounded, then so is the sequence $(e^{Dt}\cdot B(S,\varepsilon))_{t\geq 0}$. Second, there is no convenient first-order description of $\mathcal{P}(M,B(1),t)$ in which integrals do not appear. In particular we do not know whether this sequence is definable in an o-minimal expansion of $\mathbb{R}_{0}$. We will adapt to the first difference by giving substantially more complex arguments. To address the second issue, we will consider control signals over a bounded open set $\Delta$ that is different from $B(1)$. The following lemma and its corollary justify this approach.
Lemma 1.
To prove Theorem 3, we can assume without loss of generality that $M$ is of the form
(10) $\operatorname{diag}(\Lambda(a_{1},b_{1}),\ldots,\Lambda(a_{m},b_{m}),a_{m+1},\ldots,a_{l})$
where $a_{i},b_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$.
Proof.
Let $\langle M,S,T\rangle$ of the Pseudo-Orbit Safety Problem with bounded $S$. Using the algorithm for computing real Jordan forms, and permuting the blocks if necessary, compute $P,J\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ such that $M=P^{-1}JP$ and $J$ is of the form (10). For all $\varepsilon,t\geq 0$, we have that
$\mathcal{R}(M,S,B(\varepsilon),t)=P^{-1}\cdot\mathcal{R}(J,P\cdot S,P\cdot B(\varepsilon),t).$
Hence $\langle M,S,T\rangle$ is a positive instance if and only if $\langle J,P\cdot S,P\cdot T\rangle$ is. Now suppose we have computed $\hat{\varepsilon}>0$ such that $(\mathcal{R}(J,P\cdot S,B(\hat{\varepsilon}),t))_{t\geq 0}$ avoids $P\cdot T$. Then we can compute $\tilde{\varepsilon}$ such that $P\cdot B(\tilde{\varepsilon})\subseteq B(\hat{\varepsilon})$. We have that $(\mathcal{R}(M,S,B(\hat{\varepsilon}),t))_{t\geq 0}$ avoids $T$. ∎
Lemma 2.
Let $\Delta$ be a bounded and open semialgebraic set containing the origin. We can compute constants $\kappa_{1},\kappa_{2}>0$ such that for all $M\in\mathbb{R}^{d\times d}$, $S,T\subseteq\mathbb{R}^{d}$, and $\varepsilon,t\geq 0$,
$\kappa_{1}\cdot\mathcal{P}(M,S,\varepsilon\cdot\Delta,t)\subseteq\mathcal{P}(M,S,B(\varepsilon),t)\subseteq\kappa_{2}\cdot\mathcal{P}(M,S,\varepsilon\cdot\Delta,t).$
Proof.
Let $\kappa_{1},\kappa_{2}>0$ be such that $\kappa_{1}\cdot\Delta\subseteq B(1)\subseteq\kappa_{2}\cdot\Delta$. Then for all $t\geq 0$, $\mathcal{P}(M,\kappa_{1}\cdot\Delta,t)\subseteq\mathcal{P}(M,B(1),t)\subseteq\mathcal{P}(M,\kappa_{2}\cdot\Delta,t)$. It remains to recall that for every $X\subseteq\mathbb{R}^{d}$ and $\varepsilon,t\geq 0$, $\mathcal{P}(M,S,\varepsilon\cdot X,t)=\varepsilon\cdot\mathcal{P}(M,S,X,t)$. ∎
Corollary 3.
For all $M,S,T$ and $\varepsilon>0$, the sequence $(\mathcal{R}(M,S,B(\varepsilon),t))_{t\geq 0}$ avoids $T$ for all $\varepsilon>0$ if and only if $(\mathcal{R}(M,S,\varepsilon\cdot\Delta,t))_{t\geq 0}$ avoids $T$ for all $\varepsilon>0$.
Proof.
Suppose $\varepsilon>0$ is such that $(\mathcal{R}(M,S,B(\varepsilon),t))_{t\geq 0}$ avoids $T$. Then $(\mathcal{R}(M,S,\varepsilon\kappa_{1}\cdot\Delta))_{t\geq 0}$ avoids $T$. Similarly, if $(\mathcal{R}(M,S,\varepsilon\cdot\Delta))_{t\geq 0}$ avoids $T$, then so does $(\mathcal{R}(M,S,B(\varepsilon/\kappa_{2}),t))_{t\geq 0}$. ∎
We next give a few lemmas and definitions required for proving our decidability results for the Pseudo-Orbit Safety Problem, in particular the set $\Delta$ that we will be using
7.1. Preparations
In this section, fix $M$ of the form (10) and
(11) $C$ $=\operatorname{diag}(\Lambda(a_{1},0),\ldots,\Lambda(a_{m},0),a_{m+1},\ldots,a_{l}),$
(12) $D$ $=\operatorname{diag}(\Lambda(0,b_{1}),\ldots,\Lambda(0,b_{m}),\mathbf{0})$
where $M,C,D\in\mathbb{R}^{d\times d}$ and $a_{i},b_{i}\in\mathbb{R}$ for all $i$. By Lemma 6, $(C,D)$ is a decomposition of $M$. We will define a set of control inputs $\Delta$ that is specifically tailored for $M$, and study the sequence $(\mathcal{R}(M,S,\varepsilon\cdot\Delta,t))_{t\geq 0}$ via the Decomposition Method.
For $1\leq i\leq m$, let $\Delta_{i}=B((0,0),1)\subseteq\mathbb{R}^{2}$, $C_{i}=\Lambda(a_{i},0)$, and $D_{i}=\Lambda(0,b_{i})$. For $m<i\leq l$, let $\Delta_{i}=[-1,1]\subseteq\mathbb{R}$, $C_{i}=[a_{i}]$, $D_{i}=[0]$. For all $i$, let $\Lambda_{i}=C_{i}+D_{i}$. Note that $\Delta_{i}$ is semialgebraic for all $i$, and $e^{Mt}=e^{Ct}e^{Dt}$ for all $t\geq 0$. For $1\leq i\leq m$, $e^{D_{i}t}$ is a $2\times 2$ rotation matrix. Hence $e^{D_{i}}\cdot\Delta_{i}=\Delta_{i}$ for all $i$, and $e^{\Lambda_{i}t}$ preserves discs centred at the origin. Motivated by this, we define
$\Delta=\prod_{i=1}^{l}\Delta_{i}.$
Note that $\Delta\subset\mathbb{R}^{d}$ only depends on the block structure of $M$. From the construction of $\Delta$ it is immediate that for every $\varepsilon_{1},\varepsilon_{2}\geq 0$, $(\varepsilon_{1}+\varepsilon_{2})\cdot\Delta=\varepsilon_{1}\cdot\Delta+\varepsilon_{2}\cdot\Delta$. Further define
$\Omega(t)=\prod_{i=1}^{l}r_{i}(t)\cdot\Delta_{i}$
where for $1\leq i\leq l$,
[…]
The next lemma illustrates the rationale behind these definitions.
Lemma 4.
For every $\varepsilon,t\geq 0$,
$\mathcal{R}(M,S,\varepsilon\cdot\Delta,t)=e^{Mt}\cdot S+\varepsilon\cdot\Omega(t).$
Proof.
Fix $\varepsilon,t\geq 0$. By definition of $\mathcal{R}$, it suffices to show that $\mathcal{P}(M,\varepsilon\cdot\Delta,t)=\varepsilon\cdot\Omega(t)$. Because of the matching block structure of $M$ and $\Delta$,
$\mathcal{P}(M,\Delta,t)=\prod_{i=1}^{l}\mathcal{P}(\Lambda_{i},\Delta_{i},t).$
We have to show that $\mathcal{P}(\Lambda_{i},\Delta_{i},t)=r_{i}(t)\cdot\Delta_{i}$ for all $1\leq i\leq l$ and $t\geq 0$. We will do this for $1\leq i\leq m$; The case of $m<i\leq l$ is similar and easier. Fix $1\leq i\leq m$, and let $u$ be a control signal over $\Delta_{i}=B((0,0),1)$ and
[…]
We have that
$\bigl{|}e^{\Lambda_{i}(t-\tau)}u(\tau)\bigr{|}=e^{a_{i}(t-\tau)}|u(\tau)|<e^{a_{i}(t-\tau)}.$
[…]
Conversely, consider $x\in\mathbb{R}^{d}$ with $|x|<r_{i}(t)$. Define
$u(\tau)=\frac{e^{-D_{i}(t-\tau)}x}{r_{i}(t)}\in\Delta_{i}$
for $0\leq\tau\leq t$. Then
[…]
Therefore, $x\in\mathcal{P}(\Lambda_{i},\Delta_{i},t)$. ∎
Applying the decomposition method, we obtain that for every $S,T\subseteq\mathbb{R}^{d}$ and $\varepsilon,t\geq 0$,
$\mathcal{R}(M,S,\varepsilon\cdot\Delta,t)\textrm{ intersects }T\quad\Leftrightarrow\quad e^{Dt}\cdot S+\varepsilon e^{-Ct}\cdot\Omega(t)\textrm{ intersects }e^{-Ct}\cdot T.$
We next derive some properties of the sequence $(e^{-Ct}\cdot\Omega(t))_{t\geq 0}$.
Lemma 5.
For all $t\geq 0$, $e^{-Ct}\cdot\Omega(t)=\prod_{i=1}^{l}e^{-a_{i}t}\cdot r_{i}(t)\cdot\Delta_{i}$.
Proof.
For all $t\geq 0$,
$e^{-Ct}\cdot\Omega(t)=\prod_{i=1}^{l}e^{-C_{i}t}\cdot r_{i}(t)\cdot\Delta_{i}.$
It remains to recall that $e^{-C_{i}t}\cdot\Delta_{i}=e^{-a_{i}t}\cdot\Delta_{i}$ for all $i$. ∎
Corollary 6.
For every $\varepsilon_{1},\varepsilon_{2}\geq 0$,
$(\varepsilon_{1}+\varepsilon_{2})e^{-Ct}\cdot\Omega(t)=\varepsilon_{1}e^{-Ct}\cdot\Omega(t)+\varepsilon_{2}e^{-Ct}\cdot\Omega(t).$
Corollary 7.
There exists $\Phi\in\mathcal{L}_{or}$ such that for all $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$,
$x\in e^{-Ct}\cdot\Omega(t)\quad\Leftrightarrow\quad\Phi(x_{1},\ldots,x_{d},a_{1},\ldots,a_{l},t,e^{-a_{1}t},\ldots,e^{-a_{l}t}).$
Corollary 8.
There exists $\gamma>0$ such that $B(\gamma)\subseteq e^{-Ct}\cdot\Omega(t)$ for all $t\geq 1$.
Proof.
First construct $\mu>0$ such that $\widehat{r}_{i}(t)>\mu$ for all $i$ and $t\geq 1$. Then let $\gamma>0$ be such that $B(\gamma)\subseteq\prod_{i=1}^{l}\mu\cdot\Delta_{i}$. ∎
By Corollary 7, the sequence $(e^{-Ct}\cdot\Omega(t))_{t\geq 0}$ is definable in $\mathbb{R}_{\exp}$, which is o-minimal. We can now give a topological result analogous to Theorem 1. Let $\mathcal{D}$ be the closure of $(e^{Dt})_{t\geq 0}$. By Lemma 2, $\mathcal{D}$ is definable in $\mathbb{R}_{0}$.
Theorem 9.
Let $S,T\subseteq\mathbb{R}^{d}$ be definable in an o-minimal expansion $\mathbb{S}$ of $\mathbb{R}_{\exp}$, and suppose $S$ is bounded. Define
(13) $Z^{(\varepsilon)}_{t}=\{A\in\mathcal{D}\mid A\cdot S+\varepsilon e^{-Ct}\cdot\Omega(t)\textrm{ intersects }e^{-Ct}\cdot T\}$
for $\varepsilon,t\geq 0$, and
$\mu=\inf\{\varepsilon\geq 0\mid\exists N.\>\forall t\geq N\colon Z^{(\varepsilon)}_{t}\neq\emptyset\},\quad\mu\in\mathbb{R}\cup\{+\infty\}.$
Then we have the following.
  • (a)
    For every $\varepsilon>\mu$, there exist infinitely many $n\in\mathbb{N}$ such that
    $\mathcal{R}(M,S,\varepsilon\cdot\Delta,n)\>\textrm{ intersects }\>T.$
  • (b)
    For every $0\leq\varepsilon<\mu$, there exists $N\in\mathbb{N}$ such that
    $(\mathcal{R}(M,S,\varepsilon\cdot\Delta,t))_{t\geq N}\>\textrm{ avoids }\>T.$
Proof.
Observe that for all $\varepsilon,t\geq 0$,
$\mathcal{R}(M,S,\varepsilon\cdot\Delta,t)\textrm{ intersects }T\Leftrightarrow e^{Dt}\in Z^{(\varepsilon)}_{t}.$
Intuitively, $Z^{(\varepsilon)}_{t}$ is the region of $\mathcal{D}$ that must be hit by $e^{Dn}$ so that $\mathcal{R}(M,S,\varepsilon\cdot\Delta,t)$ intersects $T$. For $\varepsilon>\varepsilon_{1}$, we have that $Z_{t}^{(\varepsilon)}\supseteq Z_{t}^{(\varepsilon_{1})}$.
For every $\varepsilon>0$, the sequence $(Z^{(\varepsilon)}_{t})_{t\geq 0}$ is definable [^7] Recall that we identify $\mathbb{R}^{d\times d}$ with $\mathbb{R}^{d^{2}}$. in $\mathbb{S}$: To see this, recall that $\mathcal{D}$ is definable in $\mathbb{R}_{0}$, and the formula $\Phi\in\mathcal{L}_{\exp}$ from Corollary 7 that defines the sequence $(e^{-Ct}\cdot\Omega(t))_{t\geq 0}$. Consequently, for every $\varepsilon>0$, there exists $\varphi\in\mathcal{L}_{\mathbb{S}}$ with $k+d^{2}+2$ free variables for some $k\geq 0$, and $x_{1},\ldots,x_{k}\in\mathbb{R}$ such that for all $A\in\mathbb{R}^{d\times d}$,
$A\in Z_{t}^{(\varepsilon)}\Leftrightarrow\varphi(A_{1,1},\ldots,A_{d,d},t,\varepsilon,x_{1},\ldots,x_{k}).$
Applying Theorem 3, $(Z^{(\varepsilon)}_{t})_{t\geq 0}$ converges to a limit shape which we denote by $L_{\varepsilon}$. Moreover, because $Z_{t}^{(\varepsilon)}\subseteq\mathcal{D}$ for all $t\geq 0$ and $\mathcal{D}$ is closed, $L_{\varepsilon}\subseteq\mathcal{D}$.
To prove (a), suppose $\mu$ is finite and consider $\varepsilon>\mu$. Our goal is to prove that $e^{Dn}\in Z_{n}^{(\varepsilon)}$ for infinitely many $n\in\mathbb{N}$. Choose ${\varepsilon_{1}}\in(\mu,\varepsilon)$. Since $\varepsilon_{1}<\varepsilon$, $Z^{(\varepsilon_{1})}_{t}\subseteq Z^{(\varepsilon)}$ for all $t\geq 0$. Since ${\varepsilon_{1}}>\mu$, for all sufficiently large $t$ the set $Z^{({\varepsilon_{1}})}_{t}$ is non-empty. By Lemma 2 (b), $L_{\varepsilon_{1}}$ is non-empty. We will argue that there exists an open subset $O$ of $\mathcal{D}$ such that $Z_{t}^{(\varepsilon)}\supseteq O$ for all sufficiently large $t$.
Let $\gamma$ be as in Corollary 8, and $\delta>0$ be such that for all $A,E\in\mathcal{D}$ with $E\in B(A,\delta)$, $A\cdot S\subseteq E\cdot S+B((\varepsilon-\varepsilon_{1})\gamma)$. Such $\delta$ exists because $S$ is bounded by assumption. Consider $t\geq 0$ and $A\in Z_{t}^{(\varepsilon_{1})}$. Recall Corollary 6, and note that $B((\varepsilon-\varepsilon_{1})\gamma)\subseteq(\varepsilon-\varepsilon_{1})e^{-Ct}\cdot\Omega(t)$ by the construction of $\gamma$. For every $E\in B(A,\delta)$ we have that
$A\cdot S+\varepsilon_{1}e^{-Ct}\cdot\Omega(t)$ $\subseteq E\cdot S+B((\varepsilon-\varepsilon_{1})\gamma)+\varepsilon_{1}e^{-Ct}\cdot\Omega(t)$
$\subseteq E\cdot S+(\varepsilon-\varepsilon_{1})e^{-Ct}\cdot\Omega(t)+\varepsilon_{1}e^{-Ct}\cdot\Omega(t)$
$=E\cdot S+\varepsilon e^{-Ct}\cdot\Omega(t).$
It follows that if $E\in\mathcal{D}$, then $E\in Z^{(\varepsilon)}_{t}$. Therefore,
$Z_{t}^{(\varepsilon)}\supseteq B(Z_{t}^{(\varepsilon_{1})},\delta)\cap\mathcal{D}$
for all $t\geq 0$. Next, choose $F\in L_{\varepsilon_{1}}$. By definition of the limit shape, there exists $N\in\mathbb{N}$ such that $B(F,\delta/2)$ contains a point of $Z^{({\varepsilon_{1}})}_{t}$ for all $t\geq N$. Therefore,
$B(Z_{t}^{(\varepsilon_{1})},\delta)\cap\mathcal{D}\supseteq B(F,\delta/2)\cap\mathcal{D}$
for all $t\geq N$, and we can choose $O=B(F,\delta/2)\cap\mathcal{D}$, which is a non-empty and open subset of $\mathcal{D}$. By Lemma 2 there exist infinitely $n\in\mathbb{N}$ such that $e^{Dn}\in Z_{n}^{(\varepsilon)}$.
To prove (b), fix $0\leq\varepsilon<\mu$. Consider $U=\{t\geq 0\colon Z_{t}^{(\varepsilon)}=\emptyset\}$, which is definable in $\mathbb{S}$. By o-minimality of $\mathbb{S}$, the set $U$ is a finite union of intervals. By the construction of $\mu$ and the assumption on $\varepsilon$, the set $U$ is unbounded. Hence $[N,\infty)\subseteq U$ for some $N\in\mathbb{N}$. Finally, recall that $Z^{(\varepsilon)}_{t}=\emptyset$ implies that $\mathcal{R}(M,S,\varepsilon\cdot\Delta,t)$ does not intersect $T$. ∎
7.2. Decidability for continuous-time systems
We are now in the position to prove Theorem 3. Recall Lemma 1.
Proof of Theorem 3.
Suppose we are given $\langle M,S,T\rangle$ where $S$ is bounded and $M$ is of the form (10) with $a_{i},b_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$. Let $C,D$ be a decomposition of $M$ described in Lemma 6. By Lemma 6, equations (11) and (12) hold. Derive all constructions of Section 7.1 from $M,C,D$. Because $b_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$, by Lemma 6 the closure $\mathcal{D}$ of $(e^{Dt})_{t\geq 0}$ is semialgebraic and effectively computable. Define $Z_{t}^{(\varepsilon)}$ as in (13), and construct using Corollary 7 a formula $\psi\in\mathcal{L}_{or}$ such that for all $\varepsilon,t\geq 0$,
$Z^{(\varepsilon)}_{t}\neq\emptyset\quad\Leftrightarrow\quad\psi(\varepsilon,t,e^{-a_{1}t},\ldots,e^{-a_{l}t}).$
Consider
$E_{+}$ =$=\{\varepsilon\in\mathbb{R}\colon\exists N.\>\forall t\geq N\colon\psi(\varepsilon,t,e^{-a_{1}t},\ldots,e^{-a_{l}t})\},$
$E_{-}$ =$=\{\varepsilon\in\mathbb{R}\colon\exists N.\>\forall t\geq N\colon\lnot\psi(\varepsilon,t,e^{-a_{1}t},\ldots,e^{-a_{l}t})\}.$
By Lemma 4, both $E_{+}$ and $E_{-}$ are semialgebraic and effectively computable, and $E_{+}\cup E_{-}=\mathbb{R}$. Define
$\mu=\inf\{\varepsilon\geq 0\mid\exists N.\>\forall t\geq N\colon Z^{(\varepsilon)}_{t}\neq\emptyset\}$
and observe that $\mu=\inf(E_{+}\cap\mathbb{R}_{\geq 0})$. It follows that $\mu$ is either +$\infty$+ ∞ or belongs to $\mathbb{R}_{\geq 0}\cap\overline{\mathbb{Q}}$, and its value is effectively computable. If $\mu=0$, then $\langle M,S,T\rangle$ is a negative instance: By Theorem 9 (b), for every $\varepsilon>0$ there exist infinitely many $n\in\mathbb{N}$ such that $\mathcal{R}(M,S,\varepsilon\cdot\Delta,n)$ intersects $T$. By Lemma 2, for every $\varepsilon>0$ there exist infinitely many $n\in\mathbb{N}$ such that $\mathcal{R}(M,S,B(\varepsilon),n)$ intersects $T$.
Now suppose $\mu>0$. Because $\mu>0$, the set $E_{-}\cap\mathbb{R}_{>0}$ is non-empty. Choose $\varepsilon_{1}\in\mathbb{Q}_{>0}\cap(0,\mu)$. We have $\varepsilon_{1}\in E_{-}$ by the construction of $\mu$. Applying Lemma 4 (c), we can compute $N\in\mathbb{N}$ such that $\psi(\varepsilon_{1},t,e^{a_{1}t},\ldots,e^{a_{l}t})$ does not hold, i.e. $Z^{(\varepsilon_{1})}_{t}=\emptyset$, for all $t\geq N$. Let
$\varepsilon_{2}=\sup\{\varepsilon\geq 0\colon\mathcal{R}(M,S,\varepsilon\cdot\Delta,t)\cap T=\emptyset\textrm{ for all $t\in[0,N)$}\}$
and observe that, by Corollary 7 and Lemma 2, $\varepsilon_{2}$ can be defined without parameters in $\mathbb{R}_{\exp}$. If $\varepsilon_{2}=0$, which can be checked assuming Schanuel’s conjecture, then again $\langle M,S,T\rangle$ is a negative instance: For every $\varepsilon>0$, there exists $t\in[0,N)$ such that $\mathcal{R}(M,S,\varepsilon\cdot\Delta,t)$ intersects $T$. By Lemma 2, for every $\varepsilon>0$ there exists $t\in[0,N)$ such that $\mathcal{R}(M,S,B(\varepsilon),t)$ intersects $T$. Now suppose $\varepsilon_{2}>0$. Again using (the decision procedure for $\mathbb{R}_{\exp,\,\operatorname{bt}}$ , we can compute $\hat{\varepsilon}\in\mathbb{Q}_{>0}$ such that $\hat{\varepsilon}<\varepsilon_{1}$ and $(\mathcal{R}(M,S,\hat{\varepsilon},t))_{t\geq 0}$ avoids $T$. Applying Lemma 2, we can compute $\tilde{\varepsilon}>0$ such that $(\mathcal{R}(M,S,\tilde{\varepsilon},t))_{t\geq 0}$ avoids $T$. ∎
7.3. Decidability for discrete-time systems
We now prove decidability of the discrete-time variant. Suppose we are given an instance $\langle Q,S,T\rangle$ where $S$ is bounded and $Q$ is diagonalisable and invertible. The following is analogous to Lemma 1, and proven in the same way.
Lemma 10.
We can assume without loss of generality that
where $u_{i},v_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$ and $u_{i}\neq 0$ for all $m<i\leq l$.
Proof.
Using the algorithm for computing real Jordan forms, and permuting the blocks if necessary, compute $P,J\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$ such that $Q=P^{-1}JP$ and $J$ is of the form (14) with $u_{i},v_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$ for all $i$. Because $Q$ is invertible, $u_{i}\neq 0$ for $m<i\leq l$. For all $\varepsilon,t\geq 0$, we have that
$\mathcal{R}_{d}(Q,S,B(\varepsilon),t)=P^{-1}\cdot\mathcal{R}_{d}(J,P\cdot S,P\cdot B(\varepsilon),t)$
and hence $\mathcal{R}_{d}(Q,S,B(\varepsilon),t)\cap T=\emptyset$ if and only if $\mathcal{R}_{d}(J,P\cdot S,P\cdot B(\varepsilon),t)\cap(P\cdot T)=\emptyset$. Therefore, $\langle Q,S,T\rangle$ is a positive instance if and only if $\langle J,P\cdot S,P\cdot T\rangle$ is. Now suppose we have computed $\hat{\varepsilon}>0$ such that $(\mathcal{R}_{d}(J,P\cdot S,B(\hat{\varepsilon}),t))_{t\geq 0}$ avoids $P\cdot T$. Let $\tilde{\varepsilon}$ be such that $P\cdot B(\tilde{\varepsilon})\subseteq B(\hat{\varepsilon})$. Then $(\mathcal{R}_{d}(M,S,B(\hat{\varepsilon}),t))_{t\geq 0}$ avoids $T$. ∎
Define $a_{i}=\log(|u_{i}+\bm{i}v_{i}|)$ and $b_{i}=\operatorname{Log}((u+v\bm{i})/|u+v\bm{i}|)$. Further let
  • (a)
    $M=\log(Q)$,
  • (b)
    $C=\operatorname{diag}(\Lambda(a_{1},0),\ldots,\Lambda(a_{m},0),a_{m+1},\ldots,a_{l})$, and
  • (c)
    $D=\operatorname{diag}(\Lambda(0,b_{1}),\ldots,\Lambda(0,b_{m}),\mathbf{0})\in\mathbb{R}^{d\times d}$.
Observe that $(C,D)$ is a decomposition of $M$ by Lemma 6. By, the closure $\mathcal{D}$ of $(e^{Dt})_{t\geq 0}$ is semialgebraic and effectively computable. Moreover, $Q^{n}=e^{Mn}$ for all $n\in\mathbb{N}$. Intuitively, we will study whether $\mathcal{R}_{d}(Q,S,B(\varepsilon),n)$ intersects $T$ by considering whether $\mathcal{R}(M,S,B(\varepsilon),b)$ intersects $T$. We embed the orbit $\mathcal{O}_{d}(Q,S)$ into $\mathcal{O}(M,S)$.
Assume all constructions of Section 7.1 derived from $M$. We will need the following result.
Lemma 11.
There exist $\nu_{1},\nu_{2}\in\mathbb{Q}_{>0}$ such that for all $n\in\mathbb{N}$ and $\varepsilon\geq 0$,
$\mathcal{R}(M,S,\varepsilon\nu_{1}\cdot\Delta,n)\subseteq\mathcal{R}_{d}(Q,S,B(\varepsilon),n)\subseteq\mathcal{R}(M,S,\varepsilon\nu_{2}\cdot\Delta,n).$
Proof.
It suffices to construct $\nu_{1},\nu_{2}\in\mathbb{Q}_{>0}$ such that
$\nu_{1}\cdot\Omega(n)\subseteq\sum_{i=0}^{n-1}Q^{i}\cdot B(1)\subseteq\nu_{2}\cdot\Omega(n).$
Let $\mu_{1},\mu_{2}>0$ be such that
$\mu_{1}\cdot\Delta\subseteq B(1)\subseteq\mu_{2}\cdot\Delta$
which implies that
$\mu_{1}\cdot\sum_{i=0}^{n-1}Q^{i}\cdot\Delta\subseteq\sum_{i=0}^{n-1}Q^{i}B(1)\subseteq\mu_{2}\cdot\sum_{i=0}^{n-1}Q^{i}\cdot\Delta.$
We have that $Q^{k}\cdot\Delta=\prod_{i=1}^{l}e^{a_{i}k}\cdot B_{i}$ and hence
$\sum_{i=0}^{n-1}Q^{i}\cdot\Delta=\prod_{i=1}^{l}\widehat{r}_{i}(n)\cdot\Delta_{i}$
where $\widehat{r}_{i}(n)=\frac{e^{a_{i}n}-1}{e^{a_{i}}-1}$ if $a_{i}\neq 0$ and $\widehat{r}_{i}(n)=n$ otherwise. Let $\lambda=\max\{(e^{a_{i}}-1)/a_{i}\colon a_{i}\neq 0\}$. Then
$\frac{1}{\lambda}\cdot\Omega(n)\subseteq\sum_{i=0}^{n-1}Q^{i}\cdot\Delta\subseteq\lambda\cdot\Omega(n).$
Therefore,
$\mu_{1}.\lambda\subseteq\sum_{i=0}^{n-1}Q^{i}\cdot B(1)\subseteq\lambda\mu_{2}.$
By Corollary 7 there exists $\varphi\in\mathcal{L}_{or}$ such that $x=(x_{1},\ldots,x_{d})\in e^{-Ct}\cdot\Omega(t)$ if and only if $\varphi(x,e^{-a_{1}t},\ldots,e^{-a_{l}t})$. Compute $\mu$ as in (6). Suppose $\mu=0$. Then by Theorem 9 (a), for every $\varepsilon>0$ there exist infinitely many $n\in\mathbb{N}$ such that $e^{Mn}\cdot S+\varepsilon\cdot\Omega(n)$ intersects $T$. By Lemma 11, or every $\varepsilon>0$ there exist infinitely many $n\in\mathbb{N}$ such that $\mathcal{R}_{d}(Q,S,B(\varepsilon),n)=Q^{n}S+\sum_{i=0}^{n-1}Q^{i}\cdot B(1)$ intersects $T$. Hence $\langle Q,S,T\rangle$ is a negative instance.
Now suppose $\mu>0$. Compute $\psi,E_{+},E_{-},\varepsilon_{1},N$ as in the proof of Theorem 3. We have that for all $t\geq N$, $\mathcal{R}(M,S,B(\varepsilon_{1}),t)$ does not intersect $T$. Let
$\varepsilon_{2}=\sup\{\varepsilon\geq 0\colon\mathcal{R}_{d}(Q,S,B(\varepsilon),t)\cap T=\emptyset\textrm{ for $n=0,\ldots,N-1$}\}$
and observe that, $\varepsilon_{2}$ is real algebraic and effectively computable. If $\varepsilon_{2}=0$, then again $\langle M,S,T\rangle$ is a negative instance: For every $\varepsilon>0$, there exists $0\leq n<N$ such that $\mathcal{R}_{d}(Q,S,B(\varepsilon\nu),t)$ intersects $T$. Now suppose $\varepsilon_{2}>0$. Let $\tilde{\varepsilon}=\min\{\varepsilon_{2},\varepsilon_{1}/\}$. Then $\mathcal{R}_{d}(Q,S,B(\widetilde{\varepsilon}),t)$ does not intersect $T$ for all $n\in\mathbb{N}$, and $\langle Q,S,T\rangle$ is a positive instance.
8. Discussion
The three problems we considered in this work, namely the Robust Safety Problem, the Pseudo-Orbit Safety Problem, and the Invariant Problem, are all about how well the trajectories of all points in $S$ in the dynamical system defined by $M$ can be separated from a semialgebraic set $T$. These notions of separability are related to one another, and can be organised hierarchically. We focus continuous-time systems–the picture for discrete-time systems is essentially the same. Consider $\langle M,S,T\rangle$ where $S$ is bounded. For simplicity, assume that $S$ and $T$ are closed. When this assumption is not satisfied, we must additionally do a case analysis based on whether $\operatorname{Cl}(S)$ intersects $\operatorname{Cl}(T)$. Our separability hierarchy contains the following possibilities.
  • (i)
    There exists $\varepsilon_{1}>0$ such that $(\mathcal{R}(M,S,B(\varepsilon_{1}),t))_{t\geq 0}$ avoids $T$.
  • (ii)
    There exists $\varepsilon_{2}>0$ such that $(\mathcal{R}(M,B(S,\varepsilon_{2}),B(\varepsilon_{2}),t))_{t\geq 0}$ avoids $T$. In this setting, we allow a perturbation of the initial state as well as perturbation due to an input signal. Clearly, (ii) $\Rightarrow$ (i).
  • (iii)
    There exists $\varepsilon_{3}>0$ such that $(e^{Mt}\cdot B(S,\varepsilon_{3}))_{t\geq 0}$ avoids $T$. Observe that (ii) $\Rightarrow$ (iii).
  • (iv)
    There exists an invariant $I$ of the dynamical system given by $M$ that is definable in $\mathbb{R}_{\exp}$ and separates $S$ from $T$.
  • (v)
    The sequence $(e^{Mt}\cdot S)_{t\geq 0}$ avoids $T$. Clearly, all preceding statements imply (v).
Suppose (i). We will prove (ii). For every $\tau>0$, we can construct $\varepsilon_{\tau}<\varepsilon_{1}$ such that
$\mathcal{R}(M,B(S,\varepsilon_{\tau}),B(\varepsilon_{\tau}),t)\subseteq\mathcal{R}(M,S,B(\varepsilon_{1}),t)$
for all $\tau\geq t$. Choose sufficiently small $\tau>0$ with the property that $\mathcal{R}(M,B(S,\varepsilon_{1}),B(\varepsilon_{1}),t)$ does not intersect $T$ for all $0\leq t\leq\tau$. Such $\tau$ exists because $S,T$ are closed and disjoint. We can the define $\varepsilon_{2}=\varepsilon_{\tau}$.
We showed in Section 6.2 that (iii) $\Rightarrow$ (iv). Thus (i) $\Leftrightarrow$ (ii) $\Rightarrow$ (iii) $\Rightarrow$ (iv) $\Rightarrow$ (v). In (almagor2020cont-invariants, ), the authors show that for singleton $S$, (iv) implies existence of a semialgebraic invariant $I$ that separates $S$ from $T$. Their techniques can be adapted to show the same for all bounded $S$, without the closure assumption.
The decomposition method enables us to move from linear dynamical systems to certain compact dynamical systems governed by an isometry, and is applicable to any reachability-type problem of LDS. We give an example. Consider a discrete-time LDS given by $M\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d\times d}$, $s\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{d}$, and a semialgebraic set $S\subseteq\mathbb{R}^{d}$. Let $Z=\{n\in\mathbb{N}\colon M^{n}s\in T\}$ and $\nu=\lim_{n\to\infty}\frac{|\{0\leq k<n\colon M^{k}s\in T\}|}{n}$, called the natural density of $Z$. In (kelmendi2022computing, ), it is shown that $\nu$ is always well-defined, can be approximated to arbitrary precision, and it is decidable whether $\nu=0$. The proof is based on an analysis of exponential polynomials using Weyl’s equidistribution theorem (weyl1916gleichverteilung, ) and certain lower bounds on sums of $S$-units (evertse1984sums, ). The decomposition method can be used to give an alternative, geometric proof that does not use the latter ingredient. For simplicity, assume $M$ is invertible; The non-invertible case can be handled by taking the real Jordan form and discarding the non-invertible blocks for all sufficiently large values of $n$. Compute $C,D$ as in Lemma 7. We have that $M^{n}s\in T$ if and only if $e^{Dn}s\in e^{-Ct}\cdot T$. Let $\mathcal{D}$ be the closure of $(e^{Dt})_{t\geq 0}$, $X=\mathcal{D}\cdot s$, and $L$ be the limit shape of $(e^{-Ct}\cdot T)_{t\geq 0}$. If $L$ is disjoint from $X$, then $M^{n}s\notin T$ for all sufficiently large $n$ and hence $\nu=0$. Otherwise, the value of $\nu$ is determined by how large is $L\cap X$ in $X$. More precisely, by cell decomposition in $\mathbb{R}_{\exp}$ (vdD-geometric-categories, , Sec. 4), there exist $m>0$, $k_{1},\ldots,k_{m}\geq 0$, and sequences $(S^{(i)}_{t})_{t\geq 0}$ for $1\leq i\leq m$ all definable in $\mathbb{R}_{\exp}$ such that for all $t\geq 0$, $e^{-Ct}\cdot T=\bigcup_{i=1}^{m}S^{(i)}_{t}$ and each $S^{(i)}_{t}$ is homeomorphic to $(0,1)^{k_{i}}$. That is, no degenerate topological behaviour (e.g. infinitely many holes) arise as $e^{-Ct}\cdot T$ converges to $L$. Applying Weyl’s equidistribution theorem we can deduce that $\nu$ is exactly the measure of $L\cap X$ in the natural measure space defined on the variety $X$. Because $L$ and $X$ are both semialgebraic, $\nu$ can be approximated to arbitrary precision using tools of semialgebraic geometry. To check whether $\nu=0$, on the other hand we simply compare the dimension of $L\cap X$ to that of $X$.
Finally, we discuss the use of Schanuel’s conjecture in Theorem 1.888The situation with Theorem 3 is similar. Recall from the proof that Schanuel’s conjecture is used to determine, given $M$, $T$, bounded $S$, and $N\in\mathbb{N}$, whether there exists $\varepsilon>0$ such that $e^{Mt}\cdot B(S,\varepsilon)$ does not intersect $T$ for all $0\leq t\leq N$. We refer to this problem as the Bounded-Time Safety Problem. It can be seen that $\varepsilon$ as above exists if and only if $\bigcup_{0\leq t\leq N}e^{Mt}\cdot\operatorname{Cl}(S)$ is disjoint from $\operatorname{Cl}(T)$. We will show how to construct $D>0$, $s\in\mathbb{R}^{D}$, $A\in(\mathbb{R}\cap\overline{\mathbb{Q}})^{D\times D}$, and semialgebraic $U\subseteq\mathbb{R}^{D}$ such that for all $t\geq 0$, $e^{Mt}\cdot\operatorname{Cl}(S)$ intersects $\operatorname{Cl}(T)$ if and only if $e^{At}s\in U$999 We note that this reduction does not immediately allow us to reduce the Robust Safety Problem of the Pseudo-Orbit Safety Problem to the variants with singleton $S$. Suppose the eigenvalues of $M$ are $\lambda_{1},\ldots,\lambda_{d}$ with $\lambda_{i}=a_{i}+\bm{i}b_{i}$ and $a_{i},b_{i}\in\mathbb{R}\cap\overline{\mathbb{Q}}$. The idea is to first construct $\varphi\in\mathcal{L}_{or}$ such that for all $t\geq 0$, $e^{Mt}\cdot\operatorname{Cl}(S)\cap\operatorname{Cl}(T)$ if and only if
$\exists x_{1},\ldots,x_{d}\in\mathbb{R}\colon\varphi(t,e^{a_{1}t},\sin(b_{1}t),\cos(b_{1}t),\ldots,e^{a_{d}t},\sin(b_{d}t),\cos(b_{d}t)).$
Eliminating quantifiers, construct equivalent quantifier-free $\psi(t,e^{a_{1}t},\sin(b_{1}t),\cos(b_{1}t),\ldots,e^{a_{d}t},\sin(b_{d}t),\cos(b_{d}t))$. Recall that $\psi$ is a Boolean combination of polynomial (in)equalities. Let $A=\operatorname{diag}(\Lambda(a_{1},b_{1}),\ldots,\Lambda(a_{d},b_{d}),\Lambda(0,1))$ and $s\in\mathbb{R}^{2d+2}$ be the vector of all ones. For each inequality $p(x_{1},\ldots,x_{3d+1})\sim 0$ appearing in $\psi$, we can construct semialgebraic $V$ such that for all $t\geq 0$, $p(t,e^{a_{1}t},\sin(b_{1}t),\cos(b_{1}t),\ldots,e^{a_{d}t},\sin(b_{d}t),\cos(b_{d}t))\sim 0$ if and only if $e^{Mt}s\in V$. The set $U$ is the constructed from these semialgebraic sets based on the shape of $\psi$. Thus the Bounded-Time Safety Problem, and hence the Robust Safety Problem, reduce to the variant of the former with singleton initial sets. The Bounded-Time Safety Problem with a singleton $S$ and a hyperplane $T$ has been studied in (chonev2023zeros, ), where an unconditional algorithm in small dimensions is given. At the time of writing, all known approaches to this problem that do not depend on the dimension of the system require Schanuel’s conjecture.